### v3

# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

Abstract
Because it is difficult to precisely specify complex objectives, reinforcement learning policies are often optimized using proxy reward functions that only approximate the true goal. However, optimizing proxy rewards frequently leads to reward hacking: the optimized reward function ceases to be a good proxy and the resulting policy performs poorly with respect to the unspecified true reward. Principled solutions to reward hacking have been impeded by the lack of a good definition for the problem. To address this gap, we introduce a definition of reward hacking based on the correlation between proxy and true rewards for states and actions seen by a “base policy” that breaks down under optimization. We show that this definition captures reward hacking behavior across several realistic settings, including in reinforcement learning from human feedback (RLHF). Using our formulation, we show theoretically that regularization to the base policy can effectively prevent reward hacking. While the current practice in RLHF applies a KL penalty between action distributions for this purpose, our theory suggests regularizing the $\chi^{2}$ divergence between the policies’ occupancy measures can be more effective. We intuitively show the benefits of this type of regularization and demonstrate that it better mitigates reward hacking in practice across four realistic settings, including RLHF. Our code is available at https://github.com/cassidylaidlaw/orpo.
1 Introduction
A major challenge for the designers of goal-oriented AI systems is specifying a reward function that robustly captures their goals and values. Manually designing reward functions is difficult due to the ambiguities and complexity underlying real-world scenarios (Ibarz et al., 2018). An alternative is to learn reward functions from human data (Sadigh et al., 2017; Jeon et al., 2020), but these often fail to generalize outside the distribution of behavior seen during training (McKinney et al., 2023; Tien et al., 2023). Thus, a learned or hand-specified reward function is often just a proxy for the true reward underlying the system designer’s intent. Misalignment between the two objectives can lead to reward hacking: a learned policy performs well according to the proxy reward function but not according to the true reward function (Russell et al., 2010; Amodei et al., 2016; Pan et al., 2022; Skalse et al., 2022). A reward hacking policy’s behavior is often undesirable and can be especially catastrophic when deployed in safety-critical scenarios, such as autonomous driving (Krakovna et al., 2019; Turner et al., 2019; Knox et al., 2022). Unfortunately, reward hacking is a common phenomenon (Krakovna, 2018) and has harmful effects in the real world (Lum & Isaac, 2016; Corbett-Davies et al., 2017; Obermeyer et al., 2019; Milli et al., 2021; Franchi et al., 2023; Kleinberg et al., 2023).
The ideal solution to prevent reward hacking would be to perfectly align the specified proxy and unknown true reward; however, in many domains, this is impossible to achieve. For example, imagine trying to design a reward function for a self-driving car. It would need to capture multiple factors: the arrival time, passenger comfort, compliance with traffic laws, and various other considerations—many of which are difficult to robustly measure and would need to be carefully weighted against each other (Knox et al., 2022). In practice, reward hacking can occur even with significant reward engineering efforts. This is evident in the case of recommender systems, where hand-designed reward functions have led to adverse outcomes in terms of user experience (Stray et al., 2022).
Since proxy reward functions for complex tasks are nearly always misspecified in practice, what can be done to avoid reward hacking? There is a lack of principled solutions for preventing reward hacking, stemming from a more fundamental problem: defining reward hacking in a formal sense that captures realistic cases. In particular, it is difficult to characterize what makes a proxy reward function “good”. For example, Skalse et al. (2022) define a notion of “unhackability” but find that it can only hold if such an objective is a multiple of the true reward—a requirement that is unlikely in practice.
Figure 1: We present a new characterization of reward hacking and a method for preventing it. We define a proxy reward function as one that correlates with an unknown true reward function for state-action pairs sampled from some base policy. However, optimizing the proxy alone can lead to a breakdown in the correlation and worse true reward than the base policy. We show theoretically and empirically that optimizing the proxy with $\chi^{2}$ occupancy measure regularization to the base policy can allow outperforming the base policy under the unknown true reward.
[…]
We argue that proxies are chosen because they seem to capture the true objective under some base distribution of behavior. To formalize this, we define a proxy reward as one that correlates with the true reward function under the distribution of states and actions visited by a base policy. Then, we define a hackable proxy as one that induces a decrease in true reward compared to the base policy when optimized (Figure 1). This framework captures a wide range of intuitive reward hacking cases across realistic environments: traffic control, pandemic mitigation, blood glucose regulation, and reinforcement learning from human feedback (RLHF) (Figure 2).
Furthermore, our definition motivates a method for avoiding reward hacking during policy optimization by regularizing the optimized policy to be similar to the base policy. Specifically, we find that optimizing the proxy reward minus a regularization term provides a provable lower bound on improvement in true reward. The amount of regularization needed depends on the correlation between the proxy and true rewards; as the correlation becomes stronger, it is possible to use less regularization. In practice, RLHF already uses this form of regularization through a KL divergence penalty to the base policy (Stiennon et al., 2020; Bai et al., 2022), and Theorem 5.1 provides theoretical justification for why this works. However, our results suggest an important modification: rather than penalizing the KL divergence between the policies’ action distributions as currently done in RLHF, we should instead penalize the chi-squared ($\chi^{2}$) divergence between the policies’ occupancy measures. A policy’s occupancy measure (OM) is the distribution of states or state-action pairs seen by the policy when it interacts with its environment. Unlike action distribution (AD)-based metrics, occupancy measures capture not just the actions an agent takes, but also the states it reaches. We compare both OM to AD regularization and $\chi^{2}$ to KL divergence, presenting intuitive reasons why $\chi^{2}$ OM divergence may be a better regularization target in Figures 3 and 4.
Based on our theoretical results, we empirically investigate the benefits of using $\chi^{2}$ occupancy measure regularization to prevent reward hacking. Following previous work (Ho & Ermon, 2016), we implement OM regularization using a discriminator network that approximates the OM divergence between policies. We evaluate our approach in multiple reward hacking benchmark environments (Pan et al., 2022), comparing the effects of OM and AD regularization with both $\chi^{2}$ and KL divergence on policies optimized with hackable proxy reward functions. The results of our experiments demonstrate that training with occupancy measure regularization leads to better performance under the unseen true reward function in all of the environments, validating our theoretical results. In contrast, we find that it is difficult to tune AD regularization in some environments to both prevent reward hacking and allow meaningful improvement over the base policy. Furthermore, regularization with $\chi^{2}$ divergence tends to be more robust to the choice of regularization coefficient and often achieves higher true reward than regularizing with KL divergence.
Our main contributions can be summarized as follows:
  1. 1.
    We provide a new formal definition of reward hacking and show that it captures a number of realistic case studies.
  2. 2.
    Using our definition, we establish that optimizing a proxy reward with $\chi^{2}$ occupancy measure (OM) regularization leads to a provable improvement in the unknown true reward function.
  3. 3.
    We show how to practically implement $\chi^{2}$ OM regularization and demonstrate that it outperforms the current standard for preventing reward hacking via action distribution (AD) regularization.
2 Related Work
While there have been separate lines of work investigating reward hacking and exploring the use of occupancy measure divergences for other applications, to the best of our knowledge, we are the first to specifically study the application of occupancy measure regularization to the problem of reward hacking.
Reward hacking.  Some prior works establish theoretical models of reward hacking as a special case of Goodhart’s Law (Goodhart, 1984; Leike et al., 2018; Manheim & Garrabrant, 2019; Krakovna, 2019; Skalse et al., 2022; Ngo et al., 2023). Krakovna (2018) catalog many examples of reward hacking. Pan et al. (2022) categorize different types of reward misspecification and relate optimization power to reward hacking.
Safe reinforcement learning.  In the context of reward hacking, regularizing policies to be similar to an offline policy based on their action distribution KL divergence was proposed by Stiennon et al. (2020) and has since been widely employed in the context of optimizing LLMs using RLHF (Ouyang et al., 2022; Bai et al., 2022; Glaese et al., 2022). KL regularization for RLHF has been further studied by Vieillard et al. (2021), Gao et al. (2022), and Korbak et al. (2022). Nika et al. (2024) propose a type of occupancy measure regularization in RLHF but it is limited to deterministic MDPs, while we study general stochastic MDPs. Some alternative approaches to avoid reward hacking include quantilizers (Taylor, 2016), “mild” optimization (Taylor et al., 2020), and impact regularization (Turner et al., 2020). While constrained RL can prevent the misbehavior of agents that optimize flawed reward functions (Dalal et al., 2018; Chow et al., 2019; Zhang et al., 2020; Roy et al., 2022), it simply shifts the difficulty of designing a reward function to specifying a set of constraints and weights. Robust RL usually considers a misspecified transition model, but some work has explored misspecified reward functions (Derman et al., 2021; Gadot et al., 2024). Other proposals to address the reward specification problem attempt to infer the true reward function based on the given proxy reward function, environment context, and/or feedback from humans (Hadfield-Menell et al., 2017; Reddy et al., 2020; Lee et al., 2021). Gleave et al. (2021) have previously studied quantifying the similarity of reward functions.
Other divergences for regularization in RLHF.  Some work has explored using divergences other than KL divergence for regularization in RLHF. Go et al. (2023) investigate using other $f$-divergences for aligning language models. Wang et al. (2023a) explore generalizing direct preference optimization (DPO) (Rafailov et al., 2024), a type of RLHF that leverages offline RL, to other $f$-divergences. The most relevant work to ours in this area is Huang et al. (2024), which also proposes using $\chi^{2}$ divergence as a more theoretically grounded alternative to KL divergence for regularization in RLHF. Our work differs from theirs in several key aspects: we address reward hacking in general rather than focussing specifically on learning from preference data; we use online RL instead of their DPO-like algorithm; and, we validate our approach through experiments in realistic environments while their work is primarily theoretical.
Applications of occupancy measure regularization.  Occupancy measure regularization and optimization have been used for a variety of purposes. GAIL (Ho & Ermon, 2016) is an algorithm for robust imitation learning that aims to match the imitator’s occupancy measure to that of the demonstrator. Kang et al. (2018) combines GAIL with a reward function to efficiently explore using human data. Another line of work aims to find a policy with the highest-entropy occupancy measure for the purpose of exploring the state space (Hazan et al., 2019; Lee et al., 2020; Nedergaard & Cook, 2023). Various types of distributional regularization are used in model-based RL since learned models may not generalize out-of-distribution (Yang et al., 2022).
Offline reinforcement learning.  Our work bears a superficial resemblance to prior work in offline RL, where algorithms commonly use occupancy measure or action distribution-based regularization to constrain the learned policy within the training data distribution (Fujimoto et al., 2019; Lee et al., 2022; Mandal et al., 2023; He, 2023; Cheng et al., 2022; Rashidinejad et al., 2023; Xie et al., 2023). However, the settings are fundamentally different: while offline RL is limited by a lack of coverage in the training data, the difficulty in our setting is that the reward function is misspecified. While it might be possible to avoid reward hacking by using offline RL algorithms, we leave this to future work and focus on regularization-based approaches in the online RL setting.
Refer to captionSusceptibleExposedInfectiousRecovered
User: What is the capital of France?
Assistant: The capital of France is Paris.
Refer to caption04812-100-10-50Basepolicy(4.5, -7.6)Trafficcontrol(r=0.10)Refer to caption04812-100-10-50Rewardhacking policyTrue reward $R(s,a)$(9.1, -197.2)Refer to caption04812-100-10-50Regularizedpolicy(4.9, -4.2)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.06, -0.06)Pandemicmitigation(r=0.75)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.01, -0.14)Refer to caption-0.5-0.1-0.050-10-1-0.10(-0.05, -0.05)Refer to caption-0.2-0.10-1000-100-500(-0.04, -28)Glucosemonitoring(r=0.70)Refer to caption-0.2-0.10-1000-100-500(-0.00, -106)Refer to caption-0.2-0.10Proxy reward $\tilde{R}(s,a)$-1000-100-500(-0.01, -8)Refer to caption-3030102030(-0.0, 16.4)RLHF(r=0.40)Refer to caption-3030102030(1.4, 10.0)Refer to caption-3030102030(0.3, 17.0)Refer to caption00.20.400.050.1(0.07, 0.06)AI safetygridworld(r=0.40)Refer to caption00.20.400.050.1(0.45, 0.02)Refer to caption00.20.400.050.1(0.09, 0.09)Mean reward
Figure 2: Our definition of reward hacking successfully describes reward hacking behavior in four realistic environments and an illustrative gridworld. The top row shows the distribution of proxy and true reward values for state-action pairs sampled from a domain-appropriate base policy for each environment; in all environments, the proxy and true rewards are correlated. However, as shown in the middle row, this correlation breaks down if the proxy is optimized via RL, and the true reward achieved is lower than that of the base policy, which we characterize as reward hacking. In line with our theoretical results, the bottom row shows that RL with occupancy measure regularization to the base policy can prevent reward hacking while enabling an increase in true reward.
3 Preliminaries
To study reward hacking, we consider the setting of an infinite-horizon Markov decision process (MDP). An agent takes actions $a\in\mathcal{A}$ to transition between states $s\in\mathcal{S}$ over a series of timesteps $t=0,1,2,\ldots$. We assume that $\mathcal{S}$ and $\mathcal{A}$ are finite for simplicity, but our results can easily generalize to infinite state or action spaces. The first state $s_{0}$ is sampled from an initial distribution $\mu_{0}(s)$, and when an agent takes action $a_{t}$} in $s_{t}$} at time t{t}, the next state $s_{t+1}$ is reached at timestep $t+1$ with transition probability $p(s_{t+1}\mid s_{t},a_{t})$. The agent aims to optimize a reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and rewards are accumulated over time with discount factor $\gamma\in[0,1)$. A policy $\pi$ maps each state $s$ to a distribution over actions to take at that state $\pi(a\mid s)$. We define the (normalized) return of a policy $\pi$ under a reward function $R$ as
$J(\pi,R)=(1-\gamma)\,\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$ (1)
where $\mathbb{E}_{\pi}$ refers to the expectation under the distribution of states and actions induced by running the policy $\pi$ in the environment.
We define the state-action occupancy measure $\mu_{\pi}$ of a policy $\pi$ as the expected discounted number of times the agent will be in a particular state and take a specific action:
$\mu_{\pi}(s,a)=(1-\gamma)\,\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathds{1}\{s_{t}=s\wedge a_{t}=a\}\right].$ (2)
Intuitively, the occupancy measure represents the distribution of states and actions visited by the policy over time. Furthermore, combining (1) and (2), we can see that the return of a policy can be re-written as the expected reward for states and actions sampled from the occupancy measure:
$J(\pi,R)=\sum_{s,a\in\mathcal{S}\times\mathcal{A}}\mu_{\pi}(s,a)R(s,a)=\mathbb{E}_{\mu_{\pi}}[R(s,a)].$ (3)
[…]
Desiderata for a definition of reward hacking.  To understand why defining reward hacking is difficult, consider the case study of RLHF. In RLHF, optimizing a learned reward function over LLM outputs eventually leads to the LLM producing nonsensical responses Stiennon et al. (2020). This clearly satisfies our informal definition of “optimizing the proxy makes the true reward decrease.” However, what if we were to optimize $\tilde{R}(s,a)=-R(s,a)$? In this case, optimizing $\tilde{R}$ also makes the true reward go down, but arguably this is not reward hacking because $\tilde{R}$ was not a good “proxy” in the first place. Thus, a good definition of reward hacking should distinguish between “reasonable” proxies and reward functions that are clearly unrelated to the true reward function.
In the example of RLHF, when optimization leads to nonsensical outputs, we can clearly identify that the proxy has been “hacked.” However, what if optimizing the proxy produced mediocre but not obviously bad outputs? Even if optimizing the proxy does not lead to near-optimal true reward, we wouldn’t necessarily classify this behavior as reward hacking. Thus, a good definition of reward hacking should specify a threshold for the true reward: if the true reward falls below the threshold, reward hacking has occurred, and above it, it has not.
Prior definitions of reward hacking.  Due to these challenges, prior work has struggled to define reward hacking. Skalse et al. (2022) attempted to address our first desideratum of defining reasonable proxies by introducing a notion of an “unhackable” proxy reward. However, they found that a proxy is only unhackable if it is equivalent to the true reward up to scaling—a definition that is clearly too restrictive. The robust MDP literature has considered proxy rewards that differ from the true reward by at most some constant, e.g., $|R(s,a)-\tilde{R}(s,a)|\leq c$ for all $s,a$ (Derman et al., 2021). However, we find that in many realistic cases of reward hacking, proxies can differ from the true reward by arbitrarily large amounts in certain states, making these definitions inapplicable. As for our second desideratum of defining a true reward threshold, there is little existing literature.
4 Defining Reward Hacking
Despite the difficulties in formalizing reward hacking, we argue that both of our desiderata for a definition can be met by defining reward hacking with respect to a base policy. This approach enables both a precise definition of proxy rewards and a natural true reward threshold for determining when reward hacking is occurring.
Characterizing proxy rewards.  To find a realistic definition of a good proxy reward, consider the process by which system designers create proxies. If they are hand-specified, we argue that designers usually imagine a distribution of behavior and design a reward that captures the objective under that distribution. For example, we study a traffic control simulator (Lopez et al., 2018; Vinitsky et al., 2018; Wu et al., 2022; Pan et al., 2022) where a small number of autonomous cars help to regulate traffic among a larger population of human drivers. In this case, a designer might choose average vehicle speed as a proxy for traffic flow; this is a reasonable proxy because under the distribution of typical human driving behavior, higher speeds are associated with better traffic conditions.
We can formalize this intuition by requiring that the proxy and true rewards be correlated over states and actions sampled from a base policy:
Definition 4.1 (Correlated proxy reward).
An $r$-correlated proxy reward $\tilde{R}$ with respect to a base policy $\pi_{\text{base}}$} is one that has a correlation of $r>0$ with the true reward for state-action pairs sampled from the base policy:
$\textstyle\qquad\qquad\quad\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\left(\frac{\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}\right)\left(\frac{R(s,a)-J({\pi_{\text{base}}},R)}{\sigma_{R}}\right)\right]=r,$
$\textstyle\text{where}\quad\sigma_{\tilde{R}}^{2}=\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\big(\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})\big)^{2}\right]\quad\text{and}\quad\sigma_{R}^{2}=\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\big(R(s,a)-J({\pi_{\text{base}}},R)\big)^{2}\right]$
are the variances of proxy and true rewards under the base policy.
To consider an application of our definition, let’s return to the example of our traffic environment, where we define the true reward function as the negative total commute time for all vehicles. Using an autonomous vehicle policy based on a common model of human driving behavior as $\pi_{\text{base}}$}, we plot the distribution of true and proxy rewards in the top-left corner of Figure 2. The clear correlation that we observe between the two reward functions—higher average speed corresponding to lower commute times and vice versa—confirms that average speed is a correlated proxy reward according to Definition 4.1. This demonstrates that our formal definition successfully captures intuitive real-world proxy rewards.
Definition 4.1 offers key advantages over past formalisms. First, it avoids too strongly constraining proxy rewards by allowing the proxy and true reward functions to diverge arbitrarily at some state-action pairs when those pairs have low or zero occupancy measure under the base policy. Second, it encompasses both learned proxy rewards and hand-specified ones. In reward learning, we collect annotations (such as preference comparisons or scalar feedback) for states and actions collected from a rollout policy (such as the SFT model in RLHF). Then, a reward function is estimated via supervised learning over this dataset with an appropriate loss function. If the estimated reward function generalizes well in-distribution, then we should expect the true and learned rewards to be well-correlated under states and actions sampled from the rollout policy, satisfying Definition 4.1 (see Lemma A.8).
Choosing a threshold for reward hacking.  As discussed in Section 3, defining reward hacking requires specifying a threshold of true reward below which policy behavior is considered poor enough to be classified as reward hacking. Given that we already characterize proxy rewards with respect to a base policy, this same base policy provides a natural baseline for evaluating a policy that optimizes the proxy: if optimizing the proxy leads to worse true reward than the base policy achieves, then the system designer may be better off simply using the base policy.
Definition 4.2 (Hackable proxy reward).
Suppose $\tilde{R}$ is an $r$-correlated proxy with respect to $\pi_{\text{base}}$} (Definition 4.1). Then we say reward hacking occurs when a policy $\pi$ optimized for $\tilde{R}$ has lower true reward than the base policy $\pi_{\text{base}}$}, i.e., when $J(\pi,R)<J({\pi_{\text{base}}},R)$. If an optimal policy for $\tilde{R}$ exhibits reward hacking, then we say that the proxy reward $\tilde{R}$ is hackable:
[…]
For example, in the traffic control environment, optimizing average speed as a proxy reward leads to autonomous vehicles blocking highway on-ramps, a strategy that increases overall average speed by allowing highway traffic to move freely while forcing on-ramp cars to stop completely. However, the true reward becomes extremely low because commute times for the cars on the on-ramp are now arbitrarily long. This policy is thus reward hacking under Definition 4.2 since the true reward that it accumulates is lower than that achieved by typical human driving.
Verifying our definition experimentally.  To test whether Definition 4.2 accurately captures intuitive cases of reward hacking in practice, we consider four realistic environments and an illustrative gridworld. In addition to the aforementioned traffic simulator, we consider a pandemic mitigation and glucose monitoring environments, all of which were originally studied by Pan et al. (2022) as examples of reward hacking. PandemicSimulator (Kompella et al., 2020) uses a specialized SEIR (susceptible, exposed, infected, recovered) infection model to simulate the COVID-19 pandemic among a population; the policy controls the level of lockdown restrictions placed on the population by observing the results of testing. The proxy reward omits the political cost of certain decisions. SimGlucose (Man et al., 2014) is based on an FDA-approved simulator of Type 1 Diabetes patients in which a policy monitors glucose levels and administers insulin. The true reward captures patient health while the proxy reward prioritizes reducing the monetary cost of insulin.
We also study RLHF, in which LLMs are optimized based on a reward function learned from human preferences. Following Gao et al. (2022) and Coste et al. (2024), we use a large (more robust) reward model as the true reward function and a smaller (less robust) one as the proxy. Finally, as an interpretable and illustrative example, we include the tomato-watering gridworld from Leike et al. (2017). In this environment, a robot that waters tomatoes receives true reward depending on how many plants are watered; however, the proxy additionally rewards the robot for standing in a sprinkler where it appears all the tomatoes are watered. Besides the gridworld, all of our environments reflect complex, realistic tasks with large or infinite state state spaces. To test our definition of reward hacking, we construct a natural base policy for each of the five environments. See Appendix C for more details about the environments and base policies.
The top row of Figure 2 shows the distribution of true and proxy reward values for state-action pairs sampled from these base policies in each environment. We find that in all cases, the proxy rewards correlate with the true reward, satisfying our definition of a correlated proxy. The middle row of Figure 2 shows the result of optimizing for the proxies: for each environment, we plot the distribution of true and proxy reward values for state-action pairs reached by a policy optimized on the proxy reward via RL. In all cases, we see that the true reward drops significantly compared to the base policy. Thus, these intuitive cases of reward hacking are captured by Definition 4.2.
Figure 3: Unlike RLHF, which attempts to prevent reward hacking by regularizing action distribution (AD) divergence from the base policy, our results suggest regularizing using occupancy measure (OM) divergence is more effective. These plots of the glucose monitoring environment show the typical ADs and OMs of two policies. $\pi$ is close to $\pi_{\text{base}}$} in AD; it is more likely to give slightly less insulin due to the proxy reward’s monetary cost penalty. However, $\pi$’s optimization of the proxy also leads to a vastly different OM with typical glucose levels far outside the healthy range (dotted lines). Thus, regularizing ADs to be close to $\pi_{\text{base}}$} is not enough to prevent reward hacking. Instead, divergence between the OMs better captures the reward hacking behavior.
[…]
5 Mitigating Reward Hacking with Occupancy Measure Regularization
We now discuss how our new definition of reward hacking motivates better methods for preventing it. Ideally, we would like to be able to optimize a proxy reward function and have it translate into an improvement in true reward over the base policy. To understand how this might be possible, consider again the traffic environment. The reward hacking policy exhibits behavior which is very unlikely under the base policy: human drivers hardly ever stop indefinitely on on-ramps. That is, optimizing the proxy reward leads to out-of-distribution states and actions where the correlation that made the proxy good in the first place breaks down. Visually, this can be seen in the second row of Figure 2: the reward hacking policy usually finds state-action pairs with high proxy reward and low true reward that were very unlikely to be reached by the base policy.
Thus, one solution to prevent reward hacking could be to optimize the proxy reward while avoiding states that are unlikely under the base policy. The following theorem formalizes this idea.
Theorem 5.1.
Suppose that $\tilde{R}$ is an $r$-correlated-proxy for the true reward function $R$, and let $\sigma_{\tilde{R}}$ and $\sigma_{R}$ be defined as in Definition 4.1. Then for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$ (i.e., $\mu_{{\pi_{\text{base}}}}(s,a)=0\Rightarrow\mu_{\pi}(s,a)=0$), we have
$\frac{J(\pi,R)-J({\pi_{\text{base}}},R)}{\sigma_{R}}\geq\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right),$ (4)
where $\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)=\mathbb{E}_{\mu_{\pi}}\left[\frac{\mu_{\pi}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}-1\right]$ is the $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{base}}}$.
See Appendix A for the proof. Equation (10) gives a lower bound on how much the policy $\pi$ improves over the base policy $\pi_{\text{base}}$} in terms of the true reward, normalized by the standard deviation of the true reward under the base policy. Although this is exactly what we would like to maximize, we must instead optimize the right-hand side (RHS) of (10) since we cannot access the true reward directly. The RHS consists of two terms: one measuring the normalized improvement of the policy $\pi$’s proxy reward over the base policy and the other penalizing the divergence of $\pi$’s occupancy measure from that of $\pi_{\text{base}}$}. Optimizing the first term alone often leads to reward hacking, but by both incentivizing $\pi$ to achieve high proxy reward and stay close to the base policy, $\pi$ can achieve high true reward.
By scaling the RHS of (10) and removing terms that are constant with respect to $\pi$, we arrive at the following regularized policy optimization objective to avoid reward hacking:
$\text{maximize}\quad J(\pi,\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\qquad\text{where}\qquad\lambda=\sigma_{\tilde{R}}\sqrt{1-r^{2}}.$ (5)
The amount of regularization needed to improve on the base policy depends on the strength of correlation $r$. The higher the correlation, the lower the regularization strength $\sqrt{1-r^{2}}$. Given the prefactor of $\frac{1}{r}$ on the RHS of (10), it may appear that a lower correlation leads to a larger gain in true reward. However, Lemma A.2 shows that in fact, the lower bound decreases as a function of $r$.
While Theorem 5.1 does not guarantee that the true reward can be increased by optimizing (5), it does at least allow us to provably avoid reward hacking. In Theorem A.3 in the appendix, we show that it is difficult to guarantee an improvement in true reward in general. However, if the lower bound on the RHS of (10) can be increased above zero—which can be tested empirically—then we know that the true reward has also increased. In our experiments in Section 6, we show that in many realistic environments it is possible to increase the true reward by optimizing (5).
Comparison to KL regularization in RLHF.  Currently, RLHF calls for the application of a KL penalty between the distributions of responses generated by the optimized policy and the SFT policy (Stiennon et al., 2020; Bai et al., 2022). In our setting, we can write this as
$\text{maximize}\quad J(\pi,\tilde{R})-\lambda(1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{\text{KL}}\big(\pi(\cdot\mid s_{t})\;\|\;\mu_{\pi_{\text{base}}}(\cdot\mid s_{t})\big)\right].$ (6)
That is, in RLHF, the expected KL divergence between the action distributions of $\pi$ and $\pi_{\text{base}}$} is penalized. Action distribution divergence is easy to calculate and optimize, and training LLMs with (6) seems to work well in practice. However, unlike our objective in (5), (6) lacks theoretical guarantees, and it is unclear if it works in other environments.
Our regularized objective in (5) differs from the regularization applied in RLHF in two key ways: our approach uses occupancy measure (OM) divergence instead of action distribution (AD) divergence, and it employs $\chi^{2}$ divergence instead of KL divergence. In this section, we provide intuition for why choosing OM over AD divergence and $\chi^{2}$ over KL divergence are more effective for preventing reward hacking. Then, in Section 6, we empirically explore applying different types of regularization to prevent reward hacking in the five environments we study.
[…]
Figure 4: Our theory also suggests reward hacking can more effectively be prevented by regularizing with $\chi^{2}$ divergence instead of KL divergence. This plot illustrates how $\chi^{2}$ regularization is more effective at preventing reward hacking in RLHF. Both divergences can be written as the expectation of a function $g\big(\log\big(\mu_{\pi}(s,a)/\mu_{{\pi_{\text{base}}}}(s,a)\big)\big)$ which increases the penalty on state-action pairs based on how far the log-ratio is from zero. The $g(\cdot)$ associated with KL divergence only increases slowly for large log-ratios, so policies trained with KL divergence may produce nonsensical text. In contrast, the $g(\cdot)$ for $\chi^{2}$ divergence increases exponentially, better constraining the LLM to produce text similar to the SFT policy.
Occupancy measure vs. action distribution regularization.  While AD regularization works well for RLHF, this may be because RLHF is essentially a contextual bandit problem, meaning that $\gamma=0$; in this case, OM and AD divergence are equivalent (see Appendix A.3). However, in other cases, AD divergence may not suffice to prevent reward hacking behavior. This is because, in longer-horizon environments, a small change in action distribution at a single state can lead to a much higher probability of reaching undesirable states. Figure 3 shows an example of this in the glucose monitoring environment: policies that are close in action distribution regularization produce vastly different patient glucose levels. Occupancy measure regularization avoids this issue by directly preventing the distribution of glucose levels from differing too much from the base policy. In Theorem A.5 in the appendix, we show that in general it is impossible to lower bound the improvement in true reward using almost any form of AD regularization, in contrast to our results on OM regularization.
$\chi^{2}$ vs. KL divergence.  Compared to KL divergence, $\chi^{2}$ divergence may be more effective for preventing reward hacking because it more strongly penalizes out-of-distribution state-action pairs. To illustrate this, we can write both divergences as expectations over functions of the log-ratio of the occupancy measures, which we denote $d(s,a)$:
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{{\pi_{\text{base}}}})=\mathbb{E}_{\mu_{\pi}}\left[d(s,a)+e^{-d(s,a)}\right]\qquad\chi^{2}(\mu_{\pi}\;\|\;\mu_{{\pi_{\text{base}}}})=\mathbb{E}_{\mu_{\pi}}\left[e^{d(s,a)}+e^{-d(s,a)}\right]$
$\text{where}\qquad d(s,a)=\log\big(\mu_{\pi}(s,a)/\mu_{{\pi_{\text{base}}}}(s,a)\big).$ (7)
As $d(s,a)$ increases, the optimized policy is visiting state-action pairs that are less likely under the base policy. However, KL divergence only penalizes increases in $d(s,a)$ linearly, while $\chi^{2}$ penalizes them exponentially, resulting in stronger regularization even with a lower coefficient. Figure 4 plots the functions in (7) and shows how in practice $\chi^{2}$ divergence better prevents reward hacking in RLHF.
Environment
Traffic Pandemic Glucose RLHF AI safety
Method control mitigation monitoring gridworld
$(\times 10^{3})$ $(\times 10^{3})$
Action dist. $\chi^{2}$ -1.29 $\pm$ 0.10 -12.29 $\pm$ 0.05 -74.8 $\pm$ 11.8 16.94 $\pm$ 0.07 6.24 $\pm$ 0.09
State occupancy $\chi^{2}$ -2.18 $\pm$ 0.38 -10.68 $\pm$ 0.15 -54.7 $\pm$ 1.0 9.07 $\pm$ 0.06
State-action occupancy $\chi^{2}$ -1.15 $\pm$ 0.05 -11.17 $\pm$ 0.17 -47.6 $\pm$ 0.6 9.17 $\pm$ 0.11
Action dist. KL -1.33 $\pm$ 0.05 -12.20 $\pm$ 0.06 -73.4 $\pm$ 8.3 16.81 $\pm$ 0.27 6.33 $\pm$ 0.11
State occupancy KL -1.34 $\pm$ 22.6 -10.24 $\pm$ 0.54 -58.4 $\pm$ 3.4 7.07 $\pm$ 0.11
State-action occupancy KL -1.25 $\pm$ 0.06 -11.73 $\pm$ 0.19 -48.9 $\pm$ 0.5 6.86 $\pm$ 0.17
$\pi_{\text{base}}$} -2.28 $\pm$ 0.00 -12.26 $\pm$ 0.00 -72.6 $\pm$ 0.0 16.37 $\pm$ 0.00 5.86 $\pm$ 0.00
No regularization -57.38 $\pm$ 3.53 -29.57 $\pm$ 6.86 -599.0 $\pm$ 1.6 9.16 $\pm$ 0.80 2.35 $\pm$ 0.14
Training with true reward -0.93 $\pm$ 0.11 -2.65 $\pm$ 0.83 -43.4 $\pm$ 0.8 8.54 $\pm$ 0.12
Table 1: We compare using various types of regularization to prevent reward hacking in the five environments form Figure 2. The median true reward and standard deviation across 5 random seeds is shown for the best regularization coefficient for each type of regularization. The bottom rows show results for the baselines: the base policy $\pi_{\text{base}}$}, a policy trained on the proxy reward without regularization (exhibiting reward hacking), and a policy trained on the true reward function (impossible in practice, but included as an upper bound on performance). We find that occupancy measure regularization consistently improves on action distribution regularization, and that $\chi^{2}$ divergence often outperforms KL divergence.
6 Experiments
We now show that our theoretical results—which suggest $\chi^{2}$ occupancy measure regularization can prevent reward hacking—translate to empirical success in realistic environments.
Occupancy measure regularization in practice.  Occupancy measure regularization is more difficult to implement in practice compared to action distribution regularization. While AD regularization can be added as a loss term to deep RL algorithms like proximal policy optimization (PPO) (Schulman et al., 2017), OM divergences cannot be calculated in closed form. Instead, we follow several previous works (e.g., Ho & Ermon 2016; Kang et al. 2018) and use a discriminator network to approximate OM divergences. Specifically, in Appendix B, we show the objective in (5) can be optimized via policy gradient with an adjusted reward function that depends on a discriminator $\hat{d}_{\phi}$:
$\textstyle R^{\prime}(s,a)$ $\textstyle=\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\widehat{\chi^{2}}}}e^{\hat{d}_{\phi}(s,a)}\qquad\text{where}\qquad\widehat{\chi^{2}}=\mathbb{E}_{\mu_{\pi}}\Big[\,e^{\hat{d}_{\phi}(s_{t},a_{t})}-1\,\Big]$
$\text{and}\qquad\phi$ $=\arg\min_{\phi^{\prime}}\;\mathbb{E}_{\mu_{\pi}}\Big[\,\log(1+e^{-\hat{d}_{\phi^{\prime}}(s,a)})\,\Big]+\,\mathbb{E}_{\mu_{\pi_{\text{base}}}}\Big[\,\log(1+e^{\hat{d}_{\phi^{\prime}}(s,a)})\,\Big].$ (8)
That is, optimizing a discriminator network $\hat{d}_{\phi}$ to minimize the given loss can be used to estimate and optimize $\chi^{2}$ OM divergence. We alternately train $\hat{d}_{\phi}$ via gradient descent on (8) and the policy $\pi$ via PPO based on the adjusted reward. We call this algorithm Occupancy-Regularized Policy Optimization (ORPO). See Appendix B for a full derivation of the approximations used in ORPO and Algorithm 1 for a formal description; we also describe how to regularize based on OM KL divergence.
[…]
Figure 5: The true reward achieved by policies regularized with varying amounts of action distribution or occupancy measure regularization using $\chi^{2}$ and KL divergence. The x-axis is the regularization coefficient $\lambda$ normalized by the standard deviation of proxy rewards under the base policy. Dots indicate the median reward and the shaded area is the range over random seeds. For RLHF, AD and OM regularization are equivalent, which is why OM regularization results are not shown for that column.
Experimental setup.  In each of the five environments shown in Figure 2, we train policies with four types of regularization towards the base policy: AD KL, AD $\chi^{2}$, OM KL, and OM $\chi^{2}$. In Appendix A.5, we show that Theorem 5.1 also holds for state-only occupancy measures if the environment’s reward function does not depend on the action; thus, we experiment with regularizing based on both state-action and state-only OM divergence. For each environment and type of regularization, we test a number of regularization coefficients $\lambda$. Theorem 5.1 suggests setting $\lambda=\sigma_{\tilde{R}}\sqrt{1-r^{2}}$ for an $r$-correlated proxy, so for $\chi^{2}$ regularization we test a range of values $\lambda=c\,\sigma_{\tilde{R}}$ from $c=1$ to $10^{-2}$ ($10^{-4}$ for RLHF). Since it is less clear theoretically how to set the coefficient $\lambda$ for KL regularization, we experiment with a wider range of values.
We train each combination of $\chi^{2}$ divergence, KL divergence} $\times$ {AD, state OM, state-action OM} $\times$ $\{\lambda_{1},\lambda_{2},\dots\}$ with five random seeds and measure the resulting policies’ expected returns under the true reward. As a baseline, we train a policy without any regularization, which leads to reward hacking in all environments. We also train a policy directly on the true reward function as an upper limit for performance. In RLHF, we only consider AD regularization since it is equivalent to OM regularization for LLM chatbots (see Appendix A.3). We do not train a policy on the true reward for RLHF as we found it could be hacked with enough optimization pressure. See Appendix D for all hyperparameter and experiment details.
Results.  The results of our experiments are shown in Table 1 and Figure 5. Table 1 shows the median true reward with the best coefficient for each type of regularization. We find that OM regularization consistently outperforms AD regularization across the four non-RLHF environments. In two environments (glucose monitoring and pandemic mitigation), AD regularization fails to improve on the base policy’s true reward at all. Furthermore, $\chi^{2}$ regularization tends to perform similarly to or better than KL regularization across all environments. In RLHF in particular, $\chi^{2}$ regularization leads to a larger improvement over the base policy compared to the industry-standard KL penalty, and is more stable across seeds as well.
In Figure 5, we show the true reward achieved when training with each type of regularization across a range of $\lambda$ values. In addition to performing best with an optimal coefficient, $\chi^{2}$ and OM regularization seem to also perform well over a larger range of coefficients compared to AD regularization. In Appendix E, we present the full results of our experiments and ablations of ORPO.
7 Conclusion
We have introduced a new definition for reward hacking based on the correlation between a proxy reward function and the unknown true reward that breaks down when optimizing the proxy. Furthermore, we leveraged this definition to show theoretically and empirically that $\chi^{2}$ occupancy measure regularization can more effectively prevent reward hacking than current approaches. Our results have implications for settings, like RLHF, where RL is used to optimize complex, hard-to-specify objectives. We suggest that the heuristic KL penalty used currently should be replaced by a more principled form of regularization. While OM and AD regularization are equivalent for today’s formulation of RLHF as a contextual bandit, they will likely diverge in the near future when LLM-based agents are optimized over multi-turn conversations or with tool use (Wang et al., 2023b; Abdulhai et al., 2023; Shani et al., 2024). Thus, our results provide a principled path to continuing to ensure the safety of increasingly powerful AI systems.
Acknowledgments
We would like to thank Cam Allen, Micah Carroll, Dibya Ghosh, Katie Kang, and Sam Toyer for helpful discussions and feedback on drafts.
This work was supported by the Semiconductor Research Corporation (SRC)’s Center for the Co-Design of Cognitive Systems (COCOSYS) as well as the NSF’s Human Centered Computing (HCC) program. Cassidy Laidlaw is supported by an Open Philanthropy AI Fellowship and a National Defense Science and Engineering Graduate (NDSEG) Fellowship.
Appendix
Appendix A Proofs and Additional Theoretical Results
A.1 Proof of Theorem 5.1
See 5.1
Proof.
For simplicity of exposition, define
$Z(s,a)$ $=\frac{R(s,a)-J({\pi_{\text{base}}},R)}{\sigma_{R}}\qquad\text{and}\qquad\tilde{Z}(s,a)=\frac{\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}$.
Using (3), we can rewrite (10) as
$\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$ $\leq\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$
Then, the left hand side can be rewritten as
$\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$
$\quad=\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]+\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$
$\quad=\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right],$
since by definition $\mathbb{E}_{\mu_{\pi_{\text{base}}}}[\tilde{Z}(s,a)]=\mathbb{E}_{\mu_{\pi_{\text{base}}}}[Z(s,a)]=0$. Applying the Cauchy-Schwartz inequality to this difference gives
$\mathbb{E}_{\mu_{\pi}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]-\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]$
$\quad=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]\Big[\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a)\Big]$
$\quad=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\sqrt{\mu_{\pi_{\text{base}}}(s,a)}\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)\right]\left[\frac{\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a)}{\sqrt{\mu_{\pi_{\text{base}}}(s,a)}}\right]$
$\quad\leq\sqrt{\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{\pi_{\text{base}}}(s,a)\left[\tilde{Z}(s,a)-r\,Z(s,a)\right]^{2}\right)\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{(\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a))^{2}}{\mu_{\pi_{\text{base}}}(s,a)}\right)}$
$\quad=\sqrt{\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)^{2}\right]\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$ (9)
The expectation can be calculated as
$\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\Big(\tilde{Z}(s,a)-r\,Z(s,a)\Big)^{2}\right]$
$\quad=\text{Var}_{\mu_{\pi_{\text{base}}}}\left(\tilde{Z}(s,a)-r\,Z(s,a)\right)$
$\quad=\text{Var}_{\mu_{\pi_{\text{base}}}}\left(\tilde{Z}(s,a)\right)+r^{2}\,\text{Var}_{\mu_{\pi_{\text{base}}}}\left(Z(s,a)\right)-2r\,\text{Cov}_{\mu_{\pi_{\text{base}}}}\left(\tilde{Z}(s,a),Z(s,a)\right)$
$\quad=1-r^{2},$
using the fact that both reward functions have unit variance under the base policy and that their correlation is $r$. Plugging this into (9) gives the desired result. ∎
A.1.1 Near-Optimal Base Policies
While Theorem 5.1 shows that optimizing the proxy reward with regularization can improve on the base policy’s reward, it does not guarantee that the learned policy will be near-optimal. However, a simple corollary shows that if the base policy is near-optimal, then the learned policy will also be near-optimal.
Corollary A.1.
Suppose that $\tilde{R}$ is an $r$-correlated-proxy for the true reward function $R$, and let $\sigma_{\tilde{R}}$ and $\sigma_{R}$ be defined as in Definition 4.1. Furthermore, suppose that the base policy $\pi_{\text{base}}$} is near-optimal:
$J({\pi_{\text{base}}},R)\geq\max_{\pi^{*}}J(\pi^{*},R)-\epsilon\sigma_{R},$
for some $\epsilon>0$.
Then for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$, we have
$\frac{\max_{\pi^{*}}J(\pi^{*},R)-J(\pi,R)}{\sigma_{R}}\leq\epsilon-\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right).$ (10)
This result bounds the suboptimality gap—how close the learned policy is to optimal—in terms of the suboptimality gap of the base policy and the increase in the regularized proxy reward. The proof is straightforward and follows from Theorem 5.1.
A.1.2 Understanding the Lower Bound in Theorem 5.1
Denote by
$L(\pi)=\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right)$ (11)
the lower bound on increase in the true reward which is the RHS of (10). One surprising observation is that this lower bound seems to be increasing as the proxy becomes less correlated with the true reward. This would suggest that a less correlated proxy leads to better optimization of the true reward. However, as the following lemma shows, $L(\pi)$ is actually decreasing in $r$.
Lemma A.2.
Under the same conditions as Theorem 5.1, the lower bound $L(\pi)$ satisfies
$L(\pi)\leq\frac{1-\sqrt{1-r^{2}}}{r}\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$ (12)
This shows that the lower bound can be at most a factor of $\frac{1-\sqrt{1-r^{2}}}{r}$ times the divergence between the learned and base policies’ occupancy measures. This factor is increasing in $r$ and asymptotes to $r/2$ as $r\to 0$.
Proof.
Using the same notation as the proof of Theorem 5.1, we can rewrite the lower bound as
$L(\pi)=\frac{1}{r}\left(\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right).$
Following a similar argument to that in the proof of Theorem 5.1, we can write
$\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]$ $=\mathbb{E}_{\pi}\left[\tilde{Z}(s,a)\right]-\mathbb{E}_{\pi_{\text{base}}}\left[\tilde{Z}(s,a)\right]$
$=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\sqrt{\mu_{\pi_{\text{base}}}(s,a)}\tilde{Z}(s,a)\right]\left[\frac{\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a)}{\sqrt{\mu_{\pi_{\text{base}}}(s,a)}}\right]$
$\leq\sqrt{\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{\pi_{\text{base}}}(s,a)\tilde{Z}(s,a)^{2}\right)\left(\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{(\mu_{\pi}(s,a)-\mu_{\pi_{\text{base}}}(s,a))^{2}}{\mu_{\pi_{\text{base}}}(s,a)}\right)}$
$=\sqrt{\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\tilde{Z}(s,a)^{2}\right]\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}$
$=\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$
Combining this with the definition of $L(\pi)$ gives
$L(\pi)$ $\leq\frac{1}{r}\left(\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right)$
$=\frac{1-\sqrt{1-r^{2}}}{r}\sqrt{\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)},$
which completes the proof. ∎
A.1.3 Is the Lower Bound Optimizable?
While Theorem 5.1 shows that the increase in true reward over the base policy can be lower-bounded by the the proxy reward minus a regularization term, it is not clear if it is actually possible to increase the lower bound $L(\pi)$ as defined in (11). For example, if the base policy is already optimal with respect to the proxy reward, then clearly $L(\pi)\leq 0$. As another example, suppose it possible to improve $\pi_{\text{base}}$} with respect to both the true and proxy rewards, but only by visiting a state-action pair never visited by $\pi_{\text{base}}$}. In this case, it is also impossible to improve the lower bound in Theorem 5.1 while obeying the requirement that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$.
We prove two results relating to whether $L(\pi)$ can be increased above zero. Lemma A.3 shows that in general it is difficult to show when the lower bound can be optimized—there are general counterexamples where it cannot be positive. However, Lemma A.4 shows that there are MDPs with $r$-correlated proxies for any $r$ where the lower bound can be positive.
Furthermore, as we show in our experiments, in many realistic environments it does appear that the lower bound is optimizable. Furthermore, Theorem 5.1 still allows for safe optimization of the proxy reward: even if the it not possible to increase the lower bound above zero, optimizing $L(\pi)$ will at least prevent reward hacking.
Lemma A.3.
Fix any $r\in(0,1)$. Then there is an MDP with a true reward function $R$, a proxy reward $\tilde{R}$, and a base policy $\pi_{\text{base}}$} such that $\tilde{R}$ is an $r$-correlated proxy that can be improved upon in both true and proxy reward by a policy $\pi^{*}$:
$J(\pi^{*},R)$ $>J({\pi_{\text{base}}},R)$
[…]
$\mu_{\pi^{*}}$ $\ll\mu_{\pi_{\text{base}}}.$
However, for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$,
$L(\pi)=\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right)\leq 0.$
That is, Lemma A.3 shows that there is an MDP where there exists a policy $\pi^{*}$ that improves on the base policy in both true reward and proxy reward, but it is still not possible to increase the lower bound $L(\pi)$ above zero. This suggests that it is difficult to specify general conditions under which $L(\pi)$ can exceed zero. Thus, we rely on the empirical evidence from our experiments to show that the lower bound is often optimizable.
Proof.
We consider MDPs with discount $\gamma=0$, such that the transition probabilities are not relevant; only the initial state distribution $\mu_{0}$ and the reward functions specify the MDP. We split the analysis into two cases depending on whether $r\leq 1/2$ or $r\geq 1/2$.
Case 1: $r\leq 1/2$.
We define an MDP with two states $s_{1},s_{2}$ and two actions $a_{1},a_{2}$, with initial state distribution and rewards as follows:
$\mu_{0}(s_{1})$ $=\frac{1}{1+r}\qquad$ $R(s_{1},a_{1})$ $=\sqrt{\frac{r}{1-r}}\qquad$ $\tilde{R}(s_{1},a_{1})$ $=\sqrt{\frac{r}{1-r}}$
$R(s_{1},a_{2})$ $=-\sqrt{\frac{1-r}{r}}\qquad$ $\tilde{R}(s_{1},a_{2})$ $=0$
$\mu_{0}(s_{2})$ $=\frac{r}{1+r}\qquad$ $R(s_{2},\cdot)$ $=0\qquad$ $\tilde{R}(s_{2},\cdot)$ $=-\sqrt{\frac{1-r}{r}}$
Furthermore, we define ${\pi_{\text{base}}}(a_{1}\mid s_{1})=1-r$ and ${\pi_{\text{base}}}(a_{2}\mid s_{1})=r$, and ${\pi_{\text{base}}}(a_{1}\mid s_{2})=1$. Based on this, simple algebra shows the following facts:
$\mu_{\pi_{\text{base}}}(s_{1},a_{1})$ $=\frac{1-r}{1+r}$ $\qquad\mu_{\pi_{\text{base}}}(s_{1},a_{2})$ $=\frac{r}{1+r}$
[…]
$\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}\left[R(s,a)\tilde{R}(s,a)\right]$ $=\frac{r}{1+r}.$
Based on this, it is clear that $\tilde{R}$ is an $r$-correlated proxy. Furthermore, letting $\pi^{*}(a_{1}\mid s_{1})=1$ and $\pi^{*}(a_{1}\mid s_{2})=1$, we have
$J(\pi^{*},R)$ $=\frac{1}{1+r}\sqrt{\frac{r}{1-r}}>0=J({\pi_{\text{base}}},R)$
[…]
$\mu_{\pi^{*}}$ $\ll\mu_{\pi_{\text{base}}},$
satisfying the conditions in the lemma.
Now, consider any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$. We can calculate the lower bound $L(\pi)$ (ignoring the prefactor of $\frac{1}{r}$) as
$J(\pi,\tilde{R})\sqrt{1+r}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$ (13)
Since $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$, $\pi(a_{1}\mid s_{2})=1$, so it can only differ from $\pi_{\text{base}}$} in state $s_{1}$. Let $\delta=\pi(a_{1}\mid s_{1})-{\pi_{\text{base}}}(a_{1}\mid s_{1})$. Then, we can write the first term of (13) as
$J(\pi,\tilde{R})\sqrt{1+r}=\delta\frac{1}{1+r}\sqrt{\frac{r}{1-r}}\sqrt{1+r}=\delta\sqrt{\frac{r}{1-r^{2}}}.$
Note that the $\chi^{2}$ divergence between distributions $\mu$ and $\nu$ can be alternatively written as
$\chi^{2}(\mu\|\nu)=\sum_{s,a}\frac{(\mu(s,a)-\nu(s,a))^{2}}{\nu(s,a)}.$
Therefore, the $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{base}}}$ can be lower bounded as
$\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)\geq\frac{\left(\mu_{\pi}(s_{1},a_{1})-\mu_{\pi_{\text{base}}}(s_{1},a_{1})\right)^{2}}{\mu_{\pi_{\text{base}}}(s_{1},a_{1})}=\frac{\left(\frac{1-r+\delta}{1+r}-\frac{1-r}{1+r}\right)^{2}}{\frac{1-r}{1+r}}=\frac{\delta^{2}}{1-r^{2}}.$
This leads to the bound on (13):
$rL(\pi)$ $\leq\delta\sqrt{\frac{r}{1-r^{2}}}-\sqrt{(1-r^{2})\frac{\delta^{2}}{1-r^{2}}}=\delta\sqrt{\frac{r}{1-r^{2}}}-|\delta|.$
Clearly if $\delta\leq 0$ this is non-positive, and if $\delta>0$ it is also non-positive since $\sqrt{r/(1-r^{2})}<1$ as long as $r\leq 1/2$. Thus, the lower bound cannot be increased above zero in this case.
Case 2: $r\geq 1/2$.
In this case, we define an MDP with three states $s_{1},s_{2},s_{3}$ and two actions $a_{1},a_{2}$, with initial state distribution and rewards as follows:
$\mu_{0}(s_{1})$ $=\frac{2r^{2}-2r+1}{r^{2}-r+1}\qquad$ $R(s_{1},a_{1})$ $=\sqrt{\frac{1-r}{r}}\qquad$ $\tilde{R}(s_{1},a_{1})$ $=\sqrt{\frac{1-r}{r}}$
$R(s_{1},a_{2})$ $=-\sqrt{\frac{r}{1-r}}\qquad$ $\tilde{R}(s_{1},a_{2})$ $=0$
$\mu_{0}(s_{2})$ $=\frac{(1-r)^{2}}{r^{2}-r+1}\qquad$ $R(s_{2},\cdot)$ $=0\qquad$ $\tilde{R}(s_{2},\cdot)$ $=-\sqrt{\frac{r}{1-r}}$
$\mu_{0}(s_{3})$ $=\frac{(1-r)(2r-1)}{r^{2}-r+1}\qquad$ $R(s_{3},\cdot)$ $=-\sqrt{\frac{r}{1-r}}\qquad$ $\tilde{R}(s_{3},\cdot)$ $=-\sqrt{\frac{r}{1-r}}$
We define the base policy $\pi_{\text{base}}$} as follows:
${\pi_{\text{base}}}(a_{1}\mid s_{1})$ $=\frac{r^{2}}{2r^{2}-2r+1}\qquad$ ${\pi_{\text{base}}}(a_{2}\mid s_{1})$ $=\frac{(1-r)^{2}}{2r^{2}-2r+1}$
${\pi_{\text{base}}}(a_{1}\mid s_{2})$ $=1\qquad$ ${\pi_{\text{base}}}(a_{1}\mid s_{3})$ $=1.$
As above, we can show the following facts:
$\mu_{\pi_{\text{base}}}(s_{1},a_{1})$ $=\frac{r^{2}}{r^{2}-r+1}$ $\qquad\mu_{\pi_{\text{base}}}(s_{1},a_{2})$ $=\frac{(1-r)^{2}}{r^{2}-r+1}$
[…]
$\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}\left[R(s,a)\tilde{R}(s,a)\right]$ $=\frac{r^{2}}{r^{2}-r+1}.$
Again, this shows that $\tilde{R}$ is an $r$-correlated proxy. Letting $\pi^{*}(a_{1}\mid s_{1})=\pi^{*}(a_{1}\mid s_{2})=\pi^{*}(a_{1}\mid s_{3})=1$, we have
$J(\pi^{*},R)$ $=\frac{1-r}{r^{2}-r+1}\sqrt{\frac{1-r}{r}}>0=J({\pi_{\text{base}}},R)$
[…]
$\mu_{\pi^{*}}$ $\ll\mu_{\pi_{\text{base}}},$
satisfying the conditions in the lemma.
Now, consider any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$. We can calculate the lower bound $L(\pi)$ (again ignoring the prefactor of $\frac{1}{r}$) as
$J(\pi,\tilde{R})\sqrt{\frac{r^{2}-r+1}{r}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}.$ (14)
As in the previous case, let $\delta=\pi(a_{1}\mid s_{1})-{\pi_{\text{base}}}(a_{1}\mid s_{1})$. Then, we can write the first term of (14) as
$J(\pi,\tilde{R})\sqrt{\frac{r^{2}-r+1}{r}}=\delta\frac{2r^{2}-2r+1}{r^{2}-r+1}\sqrt{\frac{1-r}{r}}\sqrt{\frac{r^{2}-r+1}{r}}=\delta\frac{2r^{2}-2r+1}{r}\sqrt{\frac{1-r}{r^{2}-r+1}}.$
The $\chi^{2}$ divergence between $\mu_{\pi}$ and $\mu_{\pi_{\text{base}}}$ can be lower bounded as
$\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)\geq\frac{\left(\mu_{\pi}(s_{1},a_{1})-\mu_{\pi_{\text{base}}}(s_{1},a_{1})\right)^{2}}{\mu_{\pi_{\text{base}}}(s_{1},a_{1})}=\frac{\left(\frac{2r^{2}-2r+1}{r^{2}-r+1}\delta\right)^{2}}{\frac{r^{2}}{r^{2}-r+1}}=\delta^{2}\left(\frac{2r^{2}-2r+1}{r}\right)^{2}\frac{1}{r^{2}-r+1}.$
This leads to the bound on (14):
$rL(\pi)$ $\leq\delta\frac{2r^{2}-2r+1}{r}\sqrt{\frac{1-r}{r^{2}-r+1}}-\sqrt{(1-r^{2})\delta^{2}\left(\frac{2r^{2}-2r+1}{r}\right)^{2}\frac{1}{r^{2}-r+1}}$
$=\frac{2r^{2}-2r+1}{r\sqrt{r^{2}-r+1}}\left(\delta\sqrt{1-r}-|\delta|\sqrt{1-r^{2}}\right).$
The prefactor is positive, and as in the previous case, if $\delta\leq 0$ then the bound in non-positive. If $\delta>0$, then the bound is also non-positive since $\sqrt{1-r}<\sqrt{1-r^{2}}$ for any $r\in(0,1)$. Thus, the lower bound cannot be increased above zero in this case either, completing the proof. ∎
Lemma A.4.
Fix any $r\in(0,1)$. Then there is an MDP with a true reward function $R$, a proxy reward $\tilde{R}$, and a base policy $\pi_{\text{base}}$} such that $\tilde{R}$ is an $r$-correlated proxy and:
  1. 1.
    There is a policy $\pi^{*}$ such that $L(\pi^{*})>0$.
  2. 2.
    Any optimal policy with respect to $L(\cdot)$ is also an optimal policy with respect to the true reward function:
    $\arg\max_{\pi}L(\pi)\subseteq\arg\max_{\pi}J(\pi,R).$
Lemma A.4 shows that no matter how low the correlation coefficient $r$ is between the true and proxy rewards, there is at least some MDP where the lower bound $L(\pi)$ can be positive. Furthermore, in this MDP, maximizing $L(\pi)$ actually leads to an optimal policy with respect to the true reward function.
Proof.
As in the proof of Lemma A.3, we construct an MDP with discount factor $\gamma=0$ that has three states and two actions. The initial state distribution and reward functions are given by
$\mu_{0}(s_{1})$ $=\frac{1-r}{4}$ $\mu_{0}(s_{2})$ $=\frac{3+r}{8}$ $\mu_{0}(s_{3})$ $=\frac{3+r}{8}$
$R(s_{1},a_{1})=\tilde{R}(s_{1},a_{1})$ $=\sqrt{2\frac{1+r}{1-r}}$ $R(s_{2},\cdot)$ $=\sqrt{2\frac{1-r}{3+r}}$ $R(s_{3},\cdot)$ $=-\sqrt{2\frac{1-r}{3+r}}$
$R(s_{1},a_{2})=R(s_{1},a_{2})$ $=-\sqrt{2\frac{1-r}{1+r}}$ $\tilde{R}(s_{2},\cdot)$ $=-\sqrt{2\frac{1-r}{3+r}}$ $\tilde{R}(s_{3},\cdot)$ $=\sqrt{2\frac{1-r}{3+r}}.$
We define the base policy $\pi_{\text{base}}$} as follows:
${\pi_{\text{base}}}(a_{1}\mid s_{1})={\pi_{\text{base}}}(a_{2}\mid s_{1})=1/2\qquad{\pi_{\text{base}}}(a_{1}\mid s_{2})=1\qquad{\pi_{\text{base}}}(a_{1}\mid s_{3})=1.$
We can show the following facts:
[…]
$\quad=\frac{1-r}{8}\sqrt{2\frac{1+r}{1-r}}-\frac{1-r}{8}\sqrt{2\frac{1+r}{1-r}}+\frac{3+r}{8}\sqrt{2\frac{1-r}{3+r}}-\frac{3+r}{8}\sqrt{2\frac{1-r}{3+r}}$
$\quad=0$
$\sigma_{R}^{2}=\sigma_{\tilde{R}}^{2}=\frac{1-r}{8}2\frac{1+r}{1-r}+\frac{1-r}{8}2\frac{1+r}{1-r}+\frac{3+r}{8}2\frac{1-r}{3+r}+\frac{3+r}{8}2\frac{1-r}{3+r}=1$
$\mathbb{E}_{\mu_{\pi_{\text{base}}}}\Big[R(s,a)\tilde{R}(s,a)\Big]$
$\quad=\frac{1-r}{8}2\frac{1+r}{1-r}+\frac{1-r}{8}2\frac{1+r}{1-r}-\frac{3+r}{8}2\frac{1-r}{3+r}-\frac{3+r}{8}2\frac{1-r}{3+r}=r.$
Thus, $\tilde{R}$ is an $r$-correlated proxy.
Since the rewards for both actions are identical at $s_{2}$ and $s_{3}$, a policy can only differ meaningfully from $\pi_{\text{base}}$} at $s_{1}$. Define
$\pi_{\Delta}(a_{1}\mid s_{1})=\frac{1}{2}+\Delta\qquad\pi_{\Delta}(a_{1}\mid s_{2})=\pi_{\Delta}(a_{1}\mid s_{3})=1.$
as any such policy where $\Delta\in[-1/2,1/2]$. Then, we can calculate the lower bound $L(\pi_{\Delta})$ as
$L(\pi_{\Delta})$ $=\frac{1}{r}\left(J(\pi_{\Delta},\tilde{R})-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi_{\Delta}}\|\mu_{\pi_{\text{base}}}\right)}\right)$
$=\frac{1}{r}\left(2\Delta\left(\frac{1-r}{8}\right)\sqrt{2\frac{1+r}{1-r}}-\sqrt{(1-r^{2})\frac{1}{2}(1-r)\Delta^{2}}\right)$
$=\frac{1}{r}\left(\Delta\sqrt{\frac{(1+r)(1-r)}{2}}-|\Delta|\sqrt{\frac{(1-r)^{2}(1+r)}{2}}\right)$
$=\frac{1}{r}\sqrt{\frac{(1+r)(1-r)}{2}}\left(\Delta-\sqrt{1-r}|\Delta|\right).$
Since $\sqrt{1-r}<1$, clearly this is maximized at $\Delta=1/2$, where
$L(\pi_{\Delta})$ $=\frac{1}{r}\sqrt{\frac{(1+r)(1-r)}{2}}\left(\frac{1}{2}-\sqrt{1-r}\frac{1}{2}\right)>0.$
Furthermore, $\pi_{1/2}$ is an optimal policy with respect to the true reward function, since $R(s_{1},a_{1})>R(s_{1},a_{2})$ and $\pi_{1/2}(a_{1}\mid s_{1})=1$. Thus, letting $\pi^{*}=\pi_{1/2}$ completes the proof. ∎
A.2 Failure of Action Distribution Regularization
As discussed in the main text, the OM regularization method we propose differs from the AD-based regularization found in previous work on RLHF. The following theorem shows that almost any form of policy optimization with action distribution regularization cannot guarantee an improvement in true reward over the base policy.
Theorem A.5.
Fix $r\in(0,1)$. Consider a policy optimization objective regularized by any $f$-divergence between the action distributions of the learned policy and the base policy:
$\text{maximize}\qquad L^{\prime}(\pi)=J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})-g\left((1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\pi(\cdot\mid s_{t})\;\|\;{\pi_{\text{base}}}(\cdot\mid s_{t})\Big)\right]\right),$
where $f:[0,\infty)\to\mathbb{R}$ is a convex, continuous function with $f(1)=0$, $g:[0,\infty)\to[0,\infty)$ is a strictly increasing function with $g(0)=0$, and $D_{f}$ is the $f$-divergence:
$D_{f}(P\;\|\;Q)=\mathbb{E}_{x\sim Q}\left[f\left(\frac{P(x)}{Q(x)}\right)\right].$
Then there is an MDP, reward functions $R,\tilde{R}$, and base policy $\pi_{\text{base}}$} such that $\tilde{R}$ is an $r$-correlated proxy for $R$, but there is a policy $\widetilde{\pi}$ such that
$L^{\prime}(\widetilde{\pi})>0\qquad\text{but}\qquad J(\widetilde{\pi},R)<J({\pi_{\text{base}}},R).$
Theorem A.5 concerns a general form of policy optimization with action distribution regularization; it considers any $f$-divergence between action distributions and any way of scaling the $f$-divergence using the function $g$. For instnce, $g$ could incorporate linear scaling of the divergence as in the KL regularization used in previous work, square-root scaling as we use with $\chi^{2}$ divergence in Theorem 5.1, or any other scaling. The theorem shows that no matter how the regularization is formulated, it cannot guarantee improvement in true reward over the base policy; there is a policy that increases the regularized objective but decreases the true reward.
Proof.
Define the inverse $g^{-1}:[0,\infty)\to[0,\infty]$ as
$g^{-1}(x)=\sup\left\{y\in[0,\infty)\;\middle|\;g(y)\leq x\right\}.$
Since $f$ is continuous and $f(1)=0$, there must be a radius $\rho>0$ such that
$|u-1|\leq\rho\qquad\Rightarrow\qquad f(u)<\frac{2g^{-1}\left(\frac{1-r}{8}\right)}{1-r}.$
We construct an MDP with discount factor
$\gamma=\begin{cases}\max\left\{1-\frac{2g^{-1}(\frac{1-r}{8})}{(1-r)f(2)},\frac{1}{1+\rho},\frac{1}{2}\right\}&\quad f(2)>0\\ \max\left\{\frac{1}{1+\rho},\frac{1}{2}\right\}&\quad\text{otherwise.}\end{cases}$
There are four states and two actions. The initial state distribution and transition probabilities are
$\mu_{0}(s_{1})$ $=\frac{1+r}{4}\quad$ $p(s_{1}\mid s_{1},a_{1})=1$ $p(s_{1}\mid s_{1},a_{2})=1$
$\mu_{0}(s_{2})$ $=\frac{1+r}{4}\quad$ $p(s_{2}\mid s_{2},a_{1})=1$ $p(s_{2}\mid s_{2},a_{2})=1$
$\mu_{0}(s_{3})$ $=\frac{(1-r)(1+\gamma)}{4}\quad$ $p(s_{3}\mid s_{3},a_{1})=1$ $p(s_{4}\mid s_{3},a_{2})=1$
$\mu_{0}(s_{4})$ $=\frac{(1-r)(1-\gamma)}{4}\quad$ $p(s_{4}\mid s_{4},a_{1})=1$ $p(s_{4}\mid s_{4},a_{2})=1.$
The reward functions only depend on the state and are given by
$R(s_{1},\cdot)$ $=1\quad$ $\tilde{R}(s_{1},\cdot)$ $=1$
$R(s_{2},\cdot)$ $=-1\quad$ $\tilde{R}(s_{2},\cdot)$ $=-1$
$R(s_{3},\cdot)$ $=1\quad$ $\tilde{R}(s_{3},\cdot)$ $=-1$
$R(s_{4},\cdot)$ $=-1\quad$ $\tilde{R}(s_{4},\cdot)$ $=1.$
In graphical form, the MDP is as follows:
$s_{1}$$R~=1\tilde{R}=1$$s_{2}$$R~=1\tilde{R}=-1$$s_{3}$$R~=1\tilde{R}=-1$$s_{4}$$R~=1\tilde{R}=1$$\mu_{0}=\frac{1+r}{4}$$\mu_{0}=\frac{1+r}{4}$$\mu_{0}=\frac{(1-r)(1+\gamma)}{4}$$\mu_{0}=\frac{(1-r)(1-\gamma)}{4}$$a_{1},a_{2}$$a_{1},a_{2}$$a_{1}$$a_{2}$$a_{1},a_{2}$
We consider the base policy defined by
${\pi_{\text{base}}}(a_{1}\mid s_{1})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{1})$ $=0$
${\pi_{\text{base}}}(a_{1}\mid s_{2})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{2})$ $=0$
${\pi_{\text{base}}}(a_{1}\mid s_{3})$ $=\gamma\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{3})$ $=1-\gamma$
${\pi_{\text{base}}}(a_{1}\mid s_{4})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{4})$ $=0.$
To compute the occupancy measure of the base policy, we first can see that clearly,
$\mu_{{\pi_{\text{base}}}}(s_{1},a_{1})=\mu_{{\pi_{\text{base}}}}(s_{2},a_{1})=\frac{1+r}{4}.$
For $s_{3}$, the agent stays in the state until it takes action $a_{2}$, so we have
$\mu_{{\pi_{\text{base}}}}(s_{3})$ $=(1-\gamma)\mu_{0}(s_{3})\left[1+\gamma^{2}+\gamma^{4}+\cdots\right]$
$=(1-\gamma)\frac{(1-r)(1+\gamma)}{4}\frac{1}{1-\gamma^{2}}$
$=\frac{1-r}{4}.$
This also implies that $\mu_{{\pi_{\text{base}}}}(s_{4})=\frac{1-r}{4}$ since the occupancy measure sums to one. Thus, we can compute
[…]
$\sigma_{R}$ $=1\qquad$ $\sigma_{\tilde{R}}$ $=1$
$\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}[R(s,a)\tilde{R}(s,a)]$ $=r.$
This confirms that $\tilde{R}$ is an $r$-correlated proxy for $R$.
Now, we define $\widetilde{\pi}$ as
${\pi_{\text{base}}}(a_{1}\mid s_{1})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{1})$ $=0$
${\pi_{\text{base}}}(a_{1}\mid s_{2})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{2})$ $=0$
${\pi_{\text{base}}}(a_{1}\mid s_{3})$ $=2\gamma-1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{3})$ $=2(1-\gamma)$
${\pi_{\text{base}}}(a_{1}\mid s_{4})$ $=1\quad$ ${\pi_{\text{base}}}(a_{2}\mid s_{4})$ $=0.$
Note that since $\gamma\geq 1/2$ by definition, the policy is well-defined. As for $\pi_{\text{base}}$}, the occupancy measure of $\widetilde{\pi}$ at $s_{1}$ and $s_{2}$ is $\frac{1+r}{4}$. For $s_{3}$, we have
$\mu_{\widetilde{\pi}}(s_{3})$ $=(1-\gamma)\mu_{0}(s_{3})\left[1+\gamma(2\gamma-1)+\gamma^{2}(2\gamma-1)^{2}+\cdots\right]$
$=(1-\gamma)\frac{(1-r)(1+\gamma)}{4}\frac{1}{1-\gamma(2\gamma-1)}$
$=\frac{(1-r)(1+\gamma)}{4(1+2\gamma)}.$
For $s_{4}$,
$\mu_{\widetilde{\pi}}(s_{4})$ $=1-\mu_{\widetilde{\pi}}(s_{1})-\mu_{\widetilde{\pi}}(s_{2})-\mu_{\widetilde{\pi}}(s_{3})$
$=\frac{1-r}{2}-\frac{(1-r)(1+\gamma)}{4(1+2\gamma)}$
$=\frac{(1-r)(1+3\gamma)}{4(1+2\gamma)}.$
Based on this, we can compute the true reward of $\widetilde{\pi}$:
$J(\widetilde{\pi},R)$ $=\mathbb{E}_{\mu_{\widetilde{\pi}}}\left[R(s,a)\right]$
$=\frac{1+r}{4}-\frac{1+r}{4}+\frac{(1-r)(1+\gamma)}{4(1+2\gamma)}-\frac{(1-r)(1+3\gamma)}{4(1+2\gamma)}$
$=-\frac{\gamma(1-r)}{2(1+2\gamma)}$
$\leq 0=J({\pi_{\text{base}}},R).$
This verifies the claim that $J(\widetilde{\pi},R)<J({\pi_{\text{base}}},R)$.
[…]
$=\frac{\gamma(1-r)}{2(1+2\gamma)}$
$\geq\frac{1-r}{8}.$ (15)
Next, we compute the regularization term, which can be written as
$g\left((1-\gamma)\mathbb{E}_{\widetilde{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\widetilde{\pi}(\cdot\mid s_{t})\;\|\;{\pi_{\text{base}}}(\cdot\mid s_{t})\Big)\right]\right)$
$\quad=g\left(\mu(s_{3})D_{f}\Big(\widetilde{\pi}(\cdot\mid s_{3})\;\|\;{\pi_{\text{base}}}(\cdot\mid s_{3})\Big)\right)$
since $\pi_{\text{base}}$} and $\widetilde{\pi}$ only differ in state $s_{3}$. We can rewrite the above as
$=g\left(\frac{(1-r)(1+\gamma)}{4(1+2\gamma)}\left[{\pi_{\text{base}}}(a_{1}\mid s_{3})f\left(\frac{\widetilde{\pi}(a_{1}\mid s_{3})}{{\pi_{\text{base}}}(a_{1}\mid s_{3})}\right)+{\pi_{\text{base}}}(a_{2}\mid s_{3})f\left(\frac{\widetilde{\pi}(a_{2}\mid s_{3})}{{\pi_{\text{base}}}(a_{2}\mid s_{3})}\right)\right]\right)$
$=g\left(\frac{(1-r)(1+\gamma)}{4(1+2\gamma)}\left[\gamma f\left(2-\frac{1}{\gamma}\right)+(1-\gamma)f\left(2\right)\right]\right).$ (16)
Note that since $\gamma\geq\frac{1}{1+\rho}$, we have
$\frac{1}{\gamma}$ $\leq 1+\rho$
$2-\frac{1}{\gamma}$ $\geq 1-\rho$
$f\left(2-\frac{1}{\gamma}\right)$ $<\frac{2g^{-1}\left(\frac{1-r}{8}\right)}{1-r}.$
Plugging this back into (16) along with the fact that $(1-\gamma)f(2)\leq\frac{2g^{-1}(\frac{1-r}{8})}{(1-r)}$, we get
$<g\left(\frac{(1-r)(1+\gamma)}{4(1+2\gamma)}\left[\gamma\frac{2g^{-1}\left(\frac{1-r}{8}\right)}{1-r}+\frac{2g^{-1}(\frac{1-r}{8})}{1-r}\right]\right)$
$\leq g\left(\frac{1-r}{4}\times\frac{4g^{-1}\left(\frac{1-r}{8}\right)}{1-r}\right)$
$\leq\frac{1-r}{8}.$
Combining this with (15), we have
$L^{\prime}(\pi)$ $=J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})-g\left((1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{f}\Big(\pi(\cdot\mid s_{t})\;\|\;{\pi_{\text{base}}}(\cdot\mid s_{t})\Big)\right]\right)$
$>\frac{1-r}{8}-0-\frac{1-r}{8}=0,$
which completes the proof. ∎
A.3 Action Distribution and Occupancy Measure Divergences in LLMs
As noted in the main text, in the current paradigm of using RLHF to train LLMs, we can show that action distribution divergence between two policies is equivalent to occupancy measure divergence. In particular, RLHF for LLMs is usually modeled as a contextual bandit.
In our setting, a contextual bandit can be defined as an MDP with $\gamma=0$; then, the return of the policy $\pi$ under a reward function $R$ is given by
$J(\pi,R)=\mathbb{E}_{s\sim\mu_{0}(\cdot),a\sim\pi(\cdot\mid s)}[R(s,a)].$
That is, a single state is sampled from the initial state distribution $\mu_{0}$, and then a single action is sampled from the policy $\pi$ conditioned on that state. RLHF for LLMs follows this setting as a prompt is sampled from a dataset, the LLM generates a response, and the reward is calculated based on the prompt and response.
In this setting, it is simple to show that the action distribution and occupancy measure divergences are equivalent.
Lemma A.6.
Let $D_{f}(P\;\|\;Q)=\mathbb{E}_{x\sim Q}\left[f\big(P(x)/Q(x)\big)\right]$ be the $f$-divergence between two distributions $P$ and $Q$. Then, for any two policies $\pi,\pi^{\prime}$ in a contextual bandit, we have
$D_{f}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})=\mathbb{E}_{s\sim\mu_{0}(\cdot)}\left[D_{f}\Big(\pi(\cdot\mid s)\;\|\;\pi^{\prime}(\cdot\mid s)\Big)\right].$
Lemma A.6 applies to any $f$-divergence, including the KL and $\chi^{2}$ divergences we study in this paper.
Proof.
We have
$D_{f}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})$ $=\mathbb{E}_{s\sim\mu_{0}(\cdot),a\sim\pi^{\prime}(\cdot\mid s)}\left[f\left(\frac{\mu_{\pi}(s,a)}{\mu_{\pi^{\prime}}(a\mid s)}\right)\right]$
$=\mathbb{E}_{s\sim\mu_{0}(\cdot)}\left[\mathbb{E}_{a\sim\pi^{\prime}(\cdot\mid s)}\left[f\left(\frac{\mu_{0}(s)\pi(a\mid s)}{\mu_{0}(s)\pi^{\prime}(a\mid s)}\right)\right]\right]$
$=\mathbb{E}_{s\sim\mu_{0}(\cdot)}\left[D_{f}(\pi\;\|\;\pi^{\prime})\right],$
which completes the proof. ∎
A.3.1 Autoregressive Environments
While most RLHF implementations use the contextual bandit formulation above for the purposes of KL regularization, one can also model training an LLM as a sequential problem where each token generated is a separate action. This formulation is no longer a contextual bandit, but we can show that the action distribution and occupancy measure KL divergences are still equivalent!
Lemma A.7.
Suppose that an environment satisfies the following conditions:
  • It is deterministic: $\mu_{0}(s_{0})=1$ for exactly one state $s_{0}$, and for all $s_{t},a_{t}\in\mathcal{S}\times\mathcal{A}$, $p(s_{t+1}\mid s_{t},a_{t})=1$ for exactly one state $s_{t+1}$.
  • Exactly one sequence of actions leads to each state: if following $a_{0},\ldots,a_{t-1}$ leads to $s$, then no other sequence of actions (of any length) can also lead to $s$.
Then, for any policies $\pi,\pi^{\prime}$, the action distribution and occupancy measure KL divergences between them are equal (removing the $(1-\gamma)$ prefactor on the action distribution divergence):
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))\right].$
Lemma A.7 applies to LLMs since one can treat the “state” of the environment after $t$ timesteps as all the tokens generated so far $w_{0}w_{1}\dots w_{t-1}$, and the actions as the next token $w_{t}$, which is then appended to the state:
$a_{t}\sim\pi(a_{t}\mid s_{t})$ $=\pi(w_{t}\mid w_{0}w_{1}\dots w_{t-1}).$
$p(s_{t+1}\mid s_{t},a_{t})$ $=\mathbf{1}\{s_{t+1}=w_{0}w_{1}\dots w_{t-1}w_{t}\}\qquad\text{where}\quad s_{t}=w_{0}w_{1}\dots w_{t-1}\quad\text{and}\quad a_{t}=w_{t}.$
Thus, regardless of the formalism used to train an LLM via RLHF, the action distribution and occupancy measure KL are equivalent. However, the conditions of Lemma A.7 are unlikely to be met by many other MDPs. Many MDPs are stochastic, violating the first assumption. Even among deterministic MDPs, it is very uncommon that only a single action sequence can lead to each state.
Proof.
Given the assumptions about the environment, we can rewrite the log-occupancy measure of a state-action pair in terms of the sum of log action probabilties over the unique sequence of actions leading to that state. Suppose $a_{0},\ldots,a_{t-1}$ is the unique action sequence leading to $s$ and that this action sequence visits states $s_{0},\ldots,s_{t-1},s$. Then
$\log\mu_{\pi}(s,a)$ $=\log\left((1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathds{1}\{s_{t}=s\wedge a_{t}=a\}\right]\right)$
$=\log\left((1-\gamma)\gamma^{t}\mathbb{P}_{\pi}(s_{t}=s\wedge a_{t}=a)\right)$
$=\log\left((1-\gamma)\gamma^{t}\prod_{i=0}^{t}\pi(a_{i}\mid s_{i})\right)$
$=\log(1-\gamma)+t\log\gamma+\sum_{i=0}^{t}\log\pi(a_{i}\mid s_{i}).$
Using this, we can rewrite the occupancy measure KL divergence as
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})$ $=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mu_{\pi}(s,a)\log\left(\frac{\mu_{\pi}(s,a)}{\mu_{\pi^{\prime}}(s,a)}\right)$
$=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\;\sum_{a_{0},\ldots,a_{t}\in\mathcal{A}^{t+1}}\;\,\mathbb{P}_{\pi}(a_{0}\wedge\cdots\wedge a_{t})\sum_{i=0}^{t}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)$
$=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\;\sum_{a_{0},\ldots,a_{t}\in\mathcal{A}^{t+1}}\;\,\left(\prod_{j=0}^{t}\pi(a_{i}\mid s_{i})\right)\sum_{i=0}^{t}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big),$ (17)
where $s_{i}$ is the state reached by taking $a_{0},\ldots,a_{i-1}$.
We will now show inductively that
$\sum_{a_{0},\ldots,a_{t}\in\mathcal{A}^{t+1}}\;\,\left(\prod_{j=0}^{t}\pi(a_{j}\mid s_{j})\right)\sum_{i=0}^{t}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)$
$\qquad=\sum_{i=0}^{t}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i})).$ (18)
Consider first if $t=0$. Then
$\sum_{a_{0}\in\mathcal{A}}\pi(a_{0}\mid s_{0})\Big(\log\pi(a_{0}\mid s_{0})-\log\pi^{\prime}(a_{0}\mid s_{0})\Big)$
$=D_{\text{KL}}(\pi(\cdot\mid s_{0})\;\|\;\pi^{\prime}(\cdot\mid s_{0}))$
$=\mathbb{P}_{\pi}(s_{0})D_{\text{KL}}(\pi(\cdot\mid s_{0})\;\|\;\pi^{\prime}(\cdot\mid s_{0})).$
Now suppose (A.3.1) holds for t1t-1. Then for $t$ we have
$\sum_{a_{0},\ldots,a_{t}\in\mathcal{A}^{t+1}}\;\,\left(\prod_{j=0}^{t}\pi(a_{j}\mid s_{j})\right)\sum_{i=0}^{t}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)$
$=\sum_{a_{0},\ldots,a_{t-1}\in\mathcal{A}^{t}}\;\,\left(\prod_{j=0}^{t-1}\pi(a_{j}\mid s_{j})\right)\sum_{a_{t}\in\mathcal{A}}\pi(a_{t}\mid s_{t})\Bigg[\log\pi(a_{t}\mid s_{t})-\log\pi^{\prime}(a_{t}\mid s_{t})$
$\hskip 213.39566pt+\sum_{i=0}^{t-1}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)\Bigg]$
$=\sum_{a_{0},\ldots,a_{t-1}\in\mathcal{A}^{t}}\;\,\left(\prod_{j=0}^{t-1}\pi(a_{j}\mid s_{j})\right)\Bigg[D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))$
$\hskip 156.49014pt+\sum_{a_{t}\in\mathcal{A}}\pi(a_{t}\mid s_{t})\sum_{i=0}^{t-1}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)\Bigg]$
$=\sum_{a_{0},\ldots,a_{t-1}\in\mathcal{A}^{t}}\;\,\left(\prod_{j=0}^{t-1}\pi(a_{j}\mid s_{j})\right)\Bigg[D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))$
$\hskip 156.49014pt+\sum_{i=0}^{t-1}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)\Bigg]$
$=\sum_{s_{t}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{t})D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))$
$\qquad+\sum_{a_{0},\ldots,a_{t-1}\in\mathcal{A}^{t}}\;\,\left(\prod_{j=0}^{t-1}\pi(a_{j}\mid s_{j})\right)\sum_{i=0}^{t-1}\Big(\log\pi(a_{i}\mid s_{i})-\log\pi^{\prime}(a_{i}\mid s_{i})\Big)$
$\overset{\text{(i)}}{=}\sum_{s_{t}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{t})D_{\text{KL}}(\pi(\cdot\mid s_{t})\;\|\;\pi^{\prime}(\cdot\mid s_{t}))+\sum_{i=0}^{t-1}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))$
$=\sum_{i=0}^{t}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i})),$
where (i) is from the inductive hypothesis.
Now, plugging (A.3.1) into (17) gives
$D_{\text{KL}}(\mu_{\pi}\;\|\;\mu_{\pi^{\prime}})$
$=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\sum_{i=0}^{t}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))$
$=(1-\gamma)\sum_{i=0}^{\infty}\sum_{t=i}^{\infty}\gamma^{t}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))$
$=(1-\gamma)\sum_{i=0}^{\infty}\sum_{s_{i}\in\mathcal{S}}\mathbb{P}_{\pi}(s_{i})D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))\sum_{t=i}^{\infty}\gamma^{t}$
$=(1-\gamma)\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))\sum_{t=i}^{\infty}\gamma^{t}\right]$
$=(1-\gamma)\mathbb{E}_{\pi}\left[\frac{\gamma^{i}}{1-\gamma}\sum_{i=0}^{\infty}D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))\right]$
$=\mathbb{E}_{\pi}\left[\gamma^{i}\sum_{i=0}^{\infty}D_{\text{KL}}(\pi(\cdot\mid s_{i})\;\|\;\pi^{\prime}(\cdot\mid s_{i}))\right],$
which is the desired result. ∎
A.4 Learned Reward Functions are $r$-Correlated
Proxy reward functions are often learned from data like ratings or preference comparisons, including in the case of RLHF. Here, we show that a learned reward function with low mean-squared error—a common objective in supervised learning—is $r$-correlated with the true reward function.
Lemma A.8.
Let $R$ be the true reward function and $\tilde{R}$ be a learned reward function. Suppose that $\mathbb{E}_{\mu_{{\pi_{\text{base}}}}}\left[(R(s,a)-\tilde{R}(s,a))^{2}\right]\leq\epsilon\sigma_{R}^{2}$. Then, the learned reward function is an $r$-correlated proxy with $r\geq 1-\epsilon$.
The assumption in Lemma A.8 is that the mean-squared error over the occupancy measure of the base policy is small. This can be achieved, for example, by learning $\tilde{R}$ via least-squares regression over a training dataset of state-action pairs sampled from the base policy. Many results in learning theory show that this results in a small mean-squared error over the distribution the training data was sampled from, i.e., exactly the assumption in Lemma A.8 (Koltchinskii, 2006).
Proof.
Throughout the proof, all expectations, variances, and covariances are with respect to $\mu_{{\pi_{\text{base}}}}$. We can rewrite the assumption using the bias-variance decomposition as
$\mathbb{E}\left[(R(s,a)-\tilde{R}(s,a))^{2}\right]$
$\quad=\mathrm{Var}\left[R(s,a)-\tilde{R}(s,a)\right]+\left(\mathbb{E}\left[R(s,a)\right]-\mathbb{E}\left[\tilde{R}(s,a)\right]\right)^{2}$
$\quad=\mathrm{Var}\left[R(s,a)\right]+\mathrm{Var}\left[\tilde{R}(s,a)\right]-2\,\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]+\left(\mathbb{E}\left[R(s,a)\right]-\mathbb{E}\left[\tilde{R}(s,a)\right]\right)^{2}$
$\quad=\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}-2\,\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]+\left(\mathbb{E}\left[R(s,a)\right]-\mathbb{E}\left[\tilde{R}(s,a)\right]\right)^{2}$
$\quad\leq\epsilon\sigma_{R}^{2}.$
Note that $\left(\mathbb{E}\left[R(s,a)\right]-\mathbb{E}\left[\tilde{R}(s,a)\right]\right)^{2}>0$, so we can rewrite the inequality as
$2\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]\geq(1-\epsilon)\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}\geq(1-\epsilon)\left(\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}\right).$
Dividing both sides by $2\sigma_{R}\sigma_{\tilde{R}}$ gives
$\frac{\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]}{\sigma_{R}\sigma_{\tilde{R}}}$ $\geq\frac{1-\epsilon}{2}\frac{\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}}{\sigma_{R}\sigma_{\tilde{R}}}$.
By the AM-GM inequality, $\frac{\sigma_{R}^{2}+\sigma_{\tilde{R}}^{2}}{2}\geq\sigma_{R}\sigma_{\tilde{R}}$, so
$\frac{\mathrm{Cov}\left[R(s,a),\tilde{R}(s,a)\right]}{\sigma_{R}\sigma_{\tilde{R}}}\geq 1-\epsilon,$
which is the desired result. ∎
A.5 State-Only Occupancy Measures
We can also consider state-only occupancy measures, defined as
$\mu_{\pi}(s)=(1-\gamma)\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathds{1}\{s_{t}=s\}\right].$
In many environments, the reward functions only depend on the state, i.e., $R(s,a)=R(s)$ and $\tilde{R}(s,a)=\tilde{R}(s)$. In this case, Theorem 5.1 holds for state-only occupancy measures as well. The proof is identical to the proof of Theorem 5.1, but replacing expectations and sums over state-action pairs with expectations over states.
Appendix B Derivation of Occupancy-Regularized Policy Optimization
In this appendix section, we show how to derive the approximations used for Occupancy-Regularized Policy Optimization (ORPO). As a reminder, we would like to optimize
$J(\pi_{\theta},\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}.$
We can rewrite its gradient as
$\nabla_{\theta}\Big(J(\pi_{\theta},\tilde{R})-\lambda\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}\Big)$
$\quad=\nabla_{\theta}J(\pi_{\theta},\tilde{R})-\frac{\lambda\nabla_{\theta}\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}}$
$\quad=\nabla_{\theta}\left(\sum_{s,a}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)\right)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}}\nabla_{\theta}\left(\sum_{s,a}\frac{\mu_{\pi_{\theta}}(s,a)^{2}}{\mu_{{\pi_{\text{base}}}}(s,a)}-1\right)$
$\quad=\sum_{s,a}\left[\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}}\nabla_{\theta}\left(\frac{\mu_{\pi_{\theta}}(s,a)^{2}}{\mu_{{\pi_{\text{base}}}}(s,a)}-1\right)\right]$
$\quad=\sum_{s,a}\left[\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\tilde{R}(s,a)-\frac{\lambda}{2\sqrt{\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)}}\frac{2\mu_{\pi_{\theta}}(s,a)}{\mu_{{\pi_{\text{base}}}}(s,a)}\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\right]$
$\quad=\sum_{s,a}\Big(\nabla_{\theta}\mu_{\pi_{\theta}}(s,a)\Big)\left(\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\chi^{2}(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}})}}\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}\right).$
As described in the main text, policy gradient algorithms can approximate this type of gradient by using an augmented reward function
$R^{\prime}(s,a)=\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\chi^{2}(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}})}}\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}.$ (19)
However, two terms in (19) cannot be computed directly: the current $\chi^{2}$ divergence between occupancy measures, and the ratio of the occupancy measures $\mu_{\pi_{\theta}}(s,a)/\mu_{\pi_{\text{base}}}(s,a)$. We first show how to approximate the latter using a discriminator network $\hat{d}_{\phi}(s,a)$, trained to optimize
$\phi=\arg\min_{\phi}\;\mathbb{E}_{\mu_{\pi_{\theta}}}\Big[\,\log(1+e^{-\hat{d}_{\phi}(s,a)})\,\Big]\\ \textstyle+\,\mathbb{E}_{\mu_{\pi_{\text{base}}}}\Big[\,\log(1+e^{\hat{d}_{\phi}(s,a)})\,\Big].$ (20)
It is well-known that exactly optimizing the loss function in (20) gives
$\hat{d}_{\phi}(s,a)=\log\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}.$ (21)
Furthermore, the OM $\chi^{2}$ divergence is given by
$\chi^{2}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right)=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[\frac{\mu_{\pi_{\theta}}(s,a)}{\mu_{\pi_{\text{base}}}(s,a)}-1\right]=\mathbb{E}_{\mu_{\pi_{\theta}}}\left[e^{\hat{d}_{\phi}(s,a)}-1\right]\eqqcolon\widehat{\chi^{2}}$. (22)
Combining (21) and (22) shows that the augmented reward in (19) can be rewritten as
$R^{\prime}(s,a)=\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\widehat{\chi^{2}}}}e^{\hat{d}_{\phi}(s,a)}.$
Putting all the steps together, the following algorithm formalizes ORPO:
Algorithm 1 Occupancy-Regularized Policy Optimization (ORPO).
1:for iteration $i=1,\dots,I$ do
2:  Collect a set of $n$ trajectories $\mathcal{D}_{\pi}$ from $\pi_{\theta}$.
3:  Collect a set of $n$ trajectories $\mathcal{D}_{\pi_{\text{base}}}$ from $\pi_{\text{base}}$}.
4:  Optimize $\phi$ via SGD to minimize
$\mathbb{E}_{\mathcal{D}_{\pi}}\Big[\,\log(1+e^{-\hat{d}_{\phi}(s,a)})\,\Big]+\,\mathbb{E}_{\mathcal{D}_{\pi_{\text{base}}}}\Big[\,\log(1+e^{\hat{d}_{\phi}(s,a)})\,\Big]$
5:  Calculate $\widehat{\chi^{2}}=\mathbb{E}_{\mathcal{D}_{\pi}}\left[e^{\hat{d}_{\phi}(s,a)}-1\right].$
6:  Transform $\mathcal{D}_{\pi}$ to $\mathcal{D}^{\prime}_{\pi}$ by replacing the rewards with $R^{\prime}(s,a)=\tilde{R}(s,a)-\frac{\lambda}{\sqrt{\widehat{\chi^{2}}}}\left(e^{\hat{d}_{\phi}(s,a)}-1\right)$.
7:  Optimize $\theta$ via SGD to minimize the proximal policy optimization (PPO) loss $L_{\text{PPO}}\left(\mathcal{D}^{\prime}_{\pi}\right)$.
8:end for
A similar approach can also be used to optimize the proxy reward regularized by KL divergence
$J(\pi_{\theta},\tilde{R})-\lambda\,D_{\text{KL}}\left(\mu_{\pi_{\theta}}\|\mu_{\pi_{\text{base}}}\right),$
by changing the augmented reward in Line 6 of Algorithm 1 to
$R^{\prime}(s,a)=\tilde{R}(s,a)-\lambda\,\hat{d}_{\phi}(s,a).$
How accurate is the discriminator-based approximation?.  To determine whether using the discriminator results in an accurate approximation of $\chi^{2}$ and $KL$ OM divergences, we plot the output of the discriminator in the glucose environment versus the theoretically correct value in Figure 6. The results suggest that the discriminator accurately approximates the log ratio of the occupancy measures, which in turn allows for accurate approximations of the OM divergences.
[…]
Figure 6: The top plot shows the state occupancy measures of the base policy $\pi_{\text{base}}$} and a policy $\pi$ optimized with ORPO in the glucose environment. The bottom plot shows the exact log ratio of the occupancy measures $\log\mu_{\pi}(s)/\mu_{\pi_{\text{base}}}(s)$ versus the discriminator output $\hat{d}_{\phi}(s)$, which attempts to approximate it. We find the discriminator output to be a good approximation of the log ratio.
Appendix C Environment Details
Here, we discuss the details of the five reward-hacking environments we study.
C.1 Traffic Control
The traffic control environment, based on the Flow simulator (Wu et al., 2022), simulates a group of human-controlled and RL-controlled vehicles on an on-ramp attempting to merge into traffic on a highway. The true reward prioritizes a small mean commute time, while the proxy reward is the average velocity of all cars. When reward hacking, the RL controlled vehicle on the on-ramp stops indefinitely and lets cars continue forward at high speeds on the highway, which maximizes the proxy reward but increases the commute times of cars on the on-ramp infinitely. As the base policy for the traffic environment we used the Intelligent Driver Model (IDM), a standard approximation of human driving behavior (Treiber et al., 2000). In practice, base policies are often learned via imitation learning, so to simulate this we generate data from the IDM controller and train a behavioral cloning (BC) policy using the generated data.
Here, the green cars are controlled by the human driver model IDM controller, and the blue cars are controlled by RL:
This particular frame showcases reward hacking behavior. The blue RL-controlled vehicle has stopped completely on the on-ramp, blocking cars behind it. This increases the average velocity of all vehicles in the simulation, as the cars on the straightway are able to continue speeding along the road without having to wait for merging cars. However, the true reward (negative average commute time) decreases endlessly as the cars on the on-ramp wait.
C.2 Glucose Monitoring
The SimGlucose blood glucose monitoring environment is an extension of the FDA-approved glucose monitoring simulator proposed by Man et al. (2014) for Type 1 Diabetes patients (Fox et al., 2020):
The RL agent (bottom) controls the insulin administered to a simulated patient in order to maintain healthy glucose levels. The true reward is a standard measure of health risk for the patient, but the proxy reward prioritizes the monetary cost of insulin. As the safe baseline policy, we train a BC policy based on data generated by a PID controller with parameters tuned by the original designers of the simulator (Steil, 2013).
In Figure 2, we adapt a diagram from Pauley et al. (2022) to represent this environment.
C.3 Pandemic Mitigation
PandemicSimulator (Kompella et al., 2020) models a population’s infection dynamics using a modified SEIR model that they designed to represent the COVID-19 pandemic:
The RL agent chooses the level of lockdown restrictions placed on the population by observing the results of testing. The proxy reward function omits the political cost associated with certain decisions. Our base policy is trained via BC on a combination of hand-specified and real-world strategies employed by governments during the pandemic, which were also used by Kompella et al. (2020) as baselines.
C.4 RLHF
We base our RLHF environment on the work of Coste et al. (2024), who study overoptimization of LLM-based reward models. The proxy reward model we use is fine-tuned from Pythia-70M (Biderman et al., 2023) on the AlpacaFarm (Dubois et al., 2023) preference dataset. As a true reward model, we use the Llama 3 Tulu V2 8B RM from AI2 (Ivison et al., 2024). As a base policy, we use Coste et al.’s SFT policy, which was fine-tuned from Pythia-1.4B on the AlpacaFarm SFT data. We evaluate policies’ true and proxy rewards on responses sampled for a held out set of prompts.
C.5 Tomato-Watering Gridworld
The tomato environment contains a sprinkler state where the agent perceives all tomatoes as being watered and thus receives high proxy reward but no true reward. We train a base policy using the true reward function, and then add a 10% chance of taking a random action to ensure there is room to improve upon it.
The gray squares in the environment represent walls, and the white squares represent open spaces where the agent can travel:
The sprinkler state is down a narrow hallway, and on the other end a tomato is down another narrow hallway.
Appendix D Experiment Details
D.1 Non-LLM Experiments
We implement ORPO using RLLib (Liang et al., 2018) and PyTorch (Paszke et al., 2019).
Network architectures
For the pandemic, traffic, and tomato-watering environments, we use policy networks based on fully-connected networks with 2 layers of 128 units, 4 layers of 512 units, and 4 units of 512 units, respectively. The policy model for the glucose environment is a basic LSTM network with 3 layers of 64 units each. We made this choice since the observation of the environment contains continuous historical information about the patient’s blood glucose levels and previously administered insulin.
The discriminator model for all four non-LLM environments is a fully connected network with 2 layers of 256 units each. We found that the discriminator architecture did not need to be tuned to each environment.
Policy initialization
Initializing using an imitation learning policy has been shown to effectively speed up the learning process (Laidlaw et al., 2023; Uchendu et al., 2023) and is used in practice for RLHF (Stiennon et al., 2020), so we initialize our policies using the specified $\pi_{\text{base}}$} for the more realistic traffic, glucose, and pandemic environments.
Hyperparameters
Some hyperparameters for the traffic environment were tuned by Pan et al. (2022). We primarily tuned the hyperparameters listed below in order to ensure that the proxy reward would be properly optimized and reward hacking would occur without regularization. This enables to see if the various regularization methods actually succeed at preventing reward hacking.
Hyperparameter Tomato Traffic Glucose Pandemic
Training iterations 500 250 500 260
Batch size 3000 40000 100000 3860
SGD minibatch size 128 16384 1024 64
SGD epochs per iteration 8 5 4 5
Optimizer Adam Adam Adam Adam
Learning rate 1e-3 5e-5 1e-4 0.0003
Gradient clipping 0.1 None 10 10
Discount rate ($\gamma$) 0.99 0.99 0.99 0.99
GAE coefficient ($\lambda$) 0.98 0.97 0.98 0.95
Entropy coefficient (start) 0.01 0.01 0.01 0.1
Entropy coefficient (end) 0.01 0.01 0.01 0.01
Entropy schedule horizon 0 0 0 500000
KL target 0.001 0.02 1e-3 0.01
Value function loss clipping 10 10,000 100 20
Value function loss coefficient 0.1 0.5 0.0001 0.5
Share value function layers F T T T
Table 2: PPO/ORPO hyperparameters.
Hyperparameter Tomato Traffic Glucose Pandemic
Discriminator reward clipping 1000 10 1e10 0.1
Regularization coefficient ($\lambda$) Varied Varied Varied Varied
$\sigma_{\tilde{R}}$ 0.05 0.0002 0.05 0.08
Table 3: ORPO-specific hyperparameters.
ORPO details
We found that a couple of tricks were useful to ensure that ORPO remained stable. First, we clip the discriminator term added to the reward functions to a range $\delta,\delta]$, since sometimes it can blow up and cause numerical issues. Second, when estimating $\widehat{\chi^{2}}$, we use a trimmed mean (trimmed by 1% in each tail) to reduce the effect that outliers have on the estimate. These are both particularly important for $\chi^{2}$ divergence, where the output of the discriminator is exponentiated for both the reward discriminator term and for estimating $\widehat{\chi^{2}}$.
D.2 RLHF Experiments
We train LLMs via the RLHF implementation used by Coste et al. (2024), which is based on OpenAssistant and trlX (Havrilla et al., 2023). To implement $\chi^{2}$ or KL regularization, we directly add a loss term to the PPO loss:
$\chi^{2}$ divergence: $\lambda\left(\frac{\pi_{\theta}(a\mid s)}{{\pi_{\text{base}}}(a\mid s)}+\frac{{\pi_{\text{base}}}(a\mid s)}{\pi_{\theta}(a\mid s)}-2\right)$
KL divergence: $\lambda\left(\log\frac{\pi_{\theta}(a\mid s)}{{\pi_{\text{base}}}(a\mid s)}+\frac{{\pi_{\text{base}}}(a\mid s)}{\pi_{\theta}(a\mid s)}-1\right)$
where $s$ is the prompt and $a$ is the sampled response. Intuitively, both loss terms have a unique minimum when $\pi_{\theta}(a\mid s)={\pi_{\text{base}}}(a\mid s)$ and in expectation are equivalent to the correct divergence. Schulman (2020) suggests that these are particularly low-variance estimates, and we find that they work well in practice.
Appendix E Additional Experiments and Results
In this appendix, we the full results from our main experiments as well as ablations of ORPO.
E.1 Win Rates for RLHF
Besides calculating the true reward for RLHF models with the Llama 3 Tulu V2 8B RM, we also calculate the win rates for the best coefficient of KL and $\chi^{2}$ regularization. We use AlpacaEval (Dubois et al., 2023) with GPT4o-mini to compute the win rate between the RLHF policy and the SFT policy. We find that the median win rate for $\chi^{2}$ divergence is higher, and it is also more consistent across random seeds. In contrast, using KL divergence leads to reward hacking for one seed and a lower median win rate.
Divergence Coefficient Median win rate Win rate range
$\chi^{2}$ 0.0008 52.83 51.98 – 53.98
KL 0.025 51.50 11.93 – 53.75
Table 4: Win rates for RLHF-trained models using KL divergence vs. $\chi^{2}$ divergence. The median win rate and range of win rates are reported across five seeds.
E.2 Results for All Regularization Coefficients
Here, we present the results of training with AD and OM regularization using $\chi^{2}$ and KL divergence across all regularization coefficients. Each table shows the median and the standard deviation of the true rewards achieved by the learned policy across 5 random seeds.
$\chi^{2}$ divergence
Coefficient AD State OM State-action OM
0.000002 -60.79 $\pm$ 8.40 -61.04 $\pm$ 2.20 -59.11 $\pm$ 4.53
0.000004 -62.01 $\pm$ 8.58 -61.85 $\pm$ 2.22 -58.21 $\pm$ 5.52
0.00001 -60.43 $\pm$ 6.61 -59.82 $\pm$ 2.85 -54.35 $\pm$ 2.81
0.00002 -62.01 $\pm$ 4.54 -56.53 $\pm$ 10.66 -1.35 $\pm$ 30.28
0.00004 -50.84 $\pm$ 6.28 -42.24 $\pm$ 22.01 -1.15 $\pm$ 0.06
0.0001 -1.29 $\pm$ 0.12 -2.18 $\pm$ 0.42 -1.38 $\pm$ 0.29
0.0002 -1.70 $\pm$ 0.12 -2.46 $\pm$ 0.65 -1.85 $\pm$ 0.29
KL divergence
Coefficient AD State OM State-action OM
0.000001 -56.20 $\pm$ 4.02 -61.59 $\pm$ 2.72 -61.18 $\pm$ 2.87
0.0000025 -59.58 $\pm$ 4.84 -54.59 $\pm$ 3.16 -59.59 $\pm$ 2.36
0.000005 -57.24 $\pm$ 2.62 -59.03 $\pm$ 5.41 -61.24 $\pm$ 2.01
0.00001 -54.84 $\pm$ 3.15 -57.62 $\pm$ 3.13 -58.85 $\pm$ 2.73
0.000025 -55.10 $\pm$ 2.64 -59.32 $\pm$ 1.37 -56.86 $\pm$ 5.48
0.00005 -49.99 $\pm$ 4.04 -59.96 $\pm$ 1.65 -53.39 $\pm$ 23.77
0.0001 -45.72 $\pm$ 9.02 -1.34 $\pm$ 25.30 -1.25 $\pm$ 0.07
0.00025 -1.33 $\pm$ 0.05 -1.47 $\pm$ 0.20 -1.51 $\pm$ 0.10
0.0005 -1.52 $\pm$ 0.04 -1.76 $\pm$ 0.22 -1.99 $\pm$ 0.35
0.001 -1.73 $\pm$ 0.07 -1.76 $\pm$ 0.26 -2.30 $\pm$ 1.14
0.0025 -1.98 $\pm$ 0.07 -1.76 $\pm$ 0.51 -1.94 $\pm$ 0.30
0.005 -2.15 $\pm$ 0.05 -1.90 $\pm$ 0.83 -2.08 $\pm$ 0.61
0.01 -2.11 $\pm$ 0.05 -2.12 $\pm$ 1.00 -2.14 $\pm$ 0.56
Table 5: All traffic control results ($\times 10^{3}$).
$\chi^{2}$ divergence
Coefficient AD State OM State-action OM
0.0008 -17.60 $\pm$ 1.78 -12.59 $\pm$ 1.63 -27.16 $\pm$ 12.83
0.0016 -36.16 $\pm$ 3.25 -11.25 $\pm$ 10.08 -11.17 $\pm$ 0.19
0.004 -34.16 $\pm$ 12.90 -10.88 $\pm$ 0.91 -11.65 $\pm$ 2.12
0.008 -12.45 $\pm$ 0.25 -10.99 $\pm$ 3.13 -12.02 $\pm$ 0.18
0.016 -12.31 $\pm$ 0.08 -10.68 $\pm$ 0.17 -12.18 $\pm$ 0.46
0.04 -12.29 $\pm$ 0.05 -10.78 $\pm$ 0.12 -12.34 $\pm$ 0.10
0.08 -12.39 $\pm$ 0.04 -10.73 $\pm$ 0.84 -12.20 $\pm$ 0.11
KL divergence
Coefficient AD State OM State-action OM
0.00006 -21.23 $\pm$ 9.74 -33.59 $\pm$ 9.89 -30.96 $\pm$ 22.42
0.00012 -30.39 $\pm$ 22.16 -41.96 $\pm$ 12.70 -19.67 $\pm$ 6.43
0.0003 -23.10 $\pm$ 5.63 -35.29 $\pm$ 10.32 -27.56 $\pm$ 7.50
0.0006 -21.85 $\pm$ 19.03 -34.56 $\pm$ 11.97 -22.40 $\pm$ 9.84
0.0012 -25.17 $\pm$ 10.24 -31.28 $\pm$ 7.83 -31.77 $\pm$ 6.01
0.003 -23.51 $\pm$ 6.91 -35.76 $\pm$ 10.53 -23.90 $\pm$ 12.81
0.006 -12.26 $\pm$ 11.82 -58.08 $\pm$ 46.98 -29.42 $\pm$ 25.37
0.012 -12.30 $\pm$ 9.27 -10.60 $\pm$ 0.87 -11.88 $\pm$ 0.81
0.03 -12.28 $\pm$ 0.14 -11.03 $\pm$ 6.86 -11.73 $\pm$ 0.21
0.06 -12.20 $\pm$ 0.07 -10.71 $\pm$ 0.18 -12.23 $\pm$ 14.25
0.12 -12.33 $\pm$ 0.04 -10.24 $\pm$ 0.61 -12.09 $\pm$ 0.38
0.3 -12.35 $\pm$ 0.04 -11.02 $\pm$ 0.57 -12.11 $\pm$ 0.25
0.6 -12.40 $\pm$ 0.04 -10.61 $\pm$ 0.36 -12.11 $\pm$ 0.28
1.2 -12.33 $\pm$ 0.03 -10.50 $\pm$ 0.26 -12.02 $\pm$ 0.28
Table 6: All pandemic mitigation results.
$\chi^{2}$ divergence
Coefficient AD State OM State-action OM
0.0005 -580.7 $\pm$ 73.8 -484.2 $\pm$ 56.6 -164.0 $\pm$ 3.5
0.001 -94.2 $\pm$ 14.0 -263.9 $\pm$ 19.2 -127.0 $\pm$ 5.9
0.0025 -97.1 $\pm$ 8.2 -146.4 $\pm$ 18.1 -93.3 $\pm$ 9.3
0.005 -76.6 $\pm$ 10.5 -109.7 $\pm$ 10.3 -55.2 $\pm$ 0.7
0.01 -84.7 $\pm$ 8.5 -72.8 $\pm$ 8.8 -47.5 $\pm$ 0.6
0.025 -85.6 $\pm$ 7.9 -54.3 $\pm$ 1.3 -50.9 $\pm$ 2.3
0.05 -74.8 $\pm$ 13.1 -57.1 $\pm$ 3.5 -113.3 $\pm$ 32.8
KL divergence
Coefficient AD State OM State-action OM
0.00003 -598.4 $\pm$ 39.7 -604.1 $\pm$ 10.7 -583.8 $\pm$ 61.2
0.00006 -600.6 $\pm$ 12.4 -589.0 $\pm$ 260.9 -594.7 $\pm$ 3.3
0.00015 -592.3 $\pm$ 51.3 -607.9 $\pm$ 11.1 -577.1 $\pm$ 11.2
0.0003 -593.6 $\pm$ 6.0 -592.0 $\pm$ 29.3 -497.9 $\pm$ 11.2
0.0006 -590.0 $\pm$ 7.5 -593.1 $\pm$ 5.4 -364.3 $\pm$ 5.8
0.0015 -459.9 $\pm$ 114.1 -511.1 $\pm$ 21.1 -181.6 $\pm$ 7.5
0.003 -270.0 $\pm$ 39.7 -332.3 $\pm$ 41.0 -101.2 $\pm$ 5.0
0.006 -154.5 $\pm$ 5.5 -158.7 $\pm$ 28.8 -61.9 $\pm$ 5.2
0.015 -84.1 $\pm$ 6.8 -82.9 $\pm$ 5.6 -48.9 $\pm$ 0.5
0.03 -98.3 $\pm$ 8.4 -58.4 $\pm$ 3.8 -49.6 $\pm$ 1.2
0.06 -88.6 $\pm$ 12.8 -59.0 $\pm$ 7.2 -78.3 $\pm$ 10.2
0.15 -82.1 $\pm$ 11.6 -75.9 $\pm$ 4.6 -106.6 $\pm$ 19.6
0.3 -73.4 $\pm$ 9.2 -98.1 $\pm$ 16.1 -127.3 $\pm$ 24.7
0.6 -88.6 $\pm$ 5.6 -112.5 $\pm$ 16.6 -118.5 $\pm$ 10.7
Table 7: All glucose monitoring results ($\times 10^{3}$).
Coefficient AD $\chi^{2}$ AD KL
0.00008 9.20 $\pm$ 0.68 8.80 $\pm$ 2.24
0.00025 14.05 $\pm$ 3.27 9.48 $\pm$ 1.02
0.0008 16.94 $\pm$ 0.07 8.84 $\pm$ 0.42
0.0025 16.84 $\pm$ 0.08 14.22 $\pm$ 2.81
0.008 16.71 $\pm$ 0.11 12.73 $\pm$ 2.75
0.025 16.59 $\pm$ 0.11 16.81 $\pm$ 0.27
0.08 16.43 $\pm$ 0.07 16.52 $\pm$ 0.08
0.25 16.22 $\pm$ 0.10 16.33 $\pm$ 0.04
0.8 16.13 $\pm$ 0.10 16.25 $\pm$ 0.13
Table 8: All results for RLHF.
$\chi^{2}$ divergence
Coefficient AD State OM State-action OM
0.0005 2.65 $\pm$ 0.67 0.64 $\pm$ 0.19 0.53 $\pm$ 0.29
0.001 3.87 $\pm$ 0.25 0.65 $\pm$ 4.63 0.60 $\pm$ 0.16
0.0025 6.24 $\pm$ 0.10 9.06 $\pm$ 0.17 9.04 $\pm$ 4.71
0.005 6.17 $\pm$ 0.03 9.06 $\pm$ 0.11 9.16 $\pm$ 0.09
0.01 6.16 $\pm$ 0.04 9.07 $\pm$ 0.07 9.17 $\pm$ 0.12
0.025 6.19 $\pm$ 0.04 8.64 $\pm$ 0.12 8.61 $\pm$ 0.18
0.05 6.14 $\pm$ 0.00 7.95 $\pm$ 0.05 7.99 $\pm$ 0.22
KL divergence
Coefficient AD State OM State-action OM
0.0008 2.52 $\pm$ 0.18 2.32 $\pm$ 0.86 2.31 $\pm$ 0.08
0.0016 2.98 $\pm$ 0.33 2.31 $\pm$ 0.07 1.11 $\pm$ 0.89
0.004 4.59 $\pm$ 0.19 2.23 $\pm$ 0.06 2.01 $\pm$ 0.23
0.008 6.10 $\pm$ 0.15 1.30 $\pm$ 0.19 1.25 $\pm$ 0.32
0.016 6.33 $\pm$ 0.13 0.82 $\pm$ 0.22 0.84 $\pm$ 0.21
0.04 6.26 $\pm$ 0.05 1.17 $\pm$ 2.92 1.81 $\pm$ 0.31
0.08 6.26 $\pm$ 0.04 7.62 $\pm$ 0.06 7.32 $\pm$ 0.28
0.16 6.21 $\pm$ 0.05 7.20 $\pm$ 0.12 7.12 $\pm$ 0.11
0.4 6.16 $\pm$ 0.03 6.89 $\pm$ 0.14 6.84 $\pm$ 0.19
0.8 6.19 $\pm$ 0.03 7.07 $\pm$ 0.12 6.86 $\pm$ 0.19
1.6 6.14 $\pm$ 0.04 6.90 $\pm$ 0.13 6.61 $\pm$ 0.31
4 6.13 $\pm$ 0.03 6.80 $\pm$ 0.16 6.79 $\pm$ 0.12
8 6.13 $\pm$ 0.01 6.81 $\pm$ 0.28 6.80 $\pm$ 0.06
16 6.13 $\pm$ 0.00 6.94 $\pm$ 0.10 6.83 $\pm$ 0.25
Table 9: All results for the tomato-watering gridworld.
E.3 Ablations
Here, we present the results of two ablations of ORPO. For each of the non-RLHF environments, we fix the optimal coefficient for state-action OM $\chi^{2}$ regularization and modify other hyperparameters. We do not ablate the RLHF experiments because RLHF is a contextual bandit (Appendix A.3) and so it isn’t actually necessary to run ORPO for RLHF. The results are shown in Table 10 below.
Order of training policy and discriminator networks.  First, we experiment with modifying ORPO to train the discriminator after the policy. In Algorithm 1, the discriminator $\hat{d}_{\phi}$ is optimized, then the rewards are updated with the discriminator outputs, and then the policy is trained with the updated rewards. An alternative is to wait to train the discriminator until after the policy has been updated, i.e., put Line 4 after Line 7. We experimented with this and found that in in most environments there is not too much difference between the two orders, although training the discriminator first gives slightly better results. However, in the pandemic mitigation environment, we found that it training the discriminator second gave results with much higher variance. This suggests that it is best to train the discriminator before augmenting the rewards to train the policy.
Discriminator reward clipping.  Second, we experiment with modifying the discriminator reward clipping parameter $\delta$ of ORPO. We found that removing the clipping parameter entirely led to NaN errors and training could not complete, so we do not report those results. However, to test sensitivity to this parameter, we tried training with a clipping parameter $10\times$ larger and $10\times$ smaller in each environment. We found that the results did not vary by much across different clipping parameters. This suggests that ORPO is relatively robust to the hyperparameter, so it does not need to be tuned precisely.
Environment
Traffic Pandemic Glucose AI safety
Method control mitigation monitoring gridworld
$(\times 10^{3})$ $(\times 10^{3})$
Default parameters -1.15 $\pm$ 0.06 -10.68 $\pm$ 0.15. -47.5 $\pm$ 0.6 9.17 $\pm$ 0.12
Train policy before discriminator -1.24 $\pm$ 0.07 -11.51 $\pm$ 0.40 -48.2 $\pm$ 0.7 8.97 $\pm$ 0.13
Discriminator reward clipping $\times 0.1$ -1.18 $\pm$ 0.14 -10.77 $\pm$ 0.16 -47.7 $\pm$ 0.8 9.21 $\pm$ 0.18
Discriminator reward clipping $\times 10$ -1.24 $\pm$ 0.03 -10.73 $\pm$ 0.37 -47.9 $\pm$ 0.7 9.20 $\pm$ 0.14
Table 10: Results of our ablations of ORPO. We report the median and standard deviation of the true reward across five random seeds for the four non-RLHF environments. The top row shows results using each environment’s most effective form of occupancy measure regularization: state-action OM for traffic, glucose, and tomato and state OM for pandemic, with the optimal $\chi^{2}$ divergence coefficients. The other rows show ablations with the same coefficient. We find that ORPO is mostly robust to different hyperparameters but that it is probably best to train the discriminator network before the policy network.